### v1

# On the optimality of coin-betting for mean estimation

Abstract
Confidence sequences are sequences of confidence sets that adapt to incoming data while maintaining validity. Recent advances have introduced an algorithmic formulation for constructing some of the tightest confidence sequences for bounded real random variables. These approaches use a coin-betting framework, where a player sequentially bets on differences between potential mean values and observed data. This letter establishes that such coin-betting formulation is optimal among all possible algorithmic frameworks for constructing confidence sequences that build on e-variables and sequential hypothesis testing.
keywords:
mean estimation , confidence sequences , coin-betting , e-variables , sequential hypothesis testing
journal: Statistics & Probability Letters
1 Introduction
Estimating the mean of a random variable from empirical observations is a classical problem in statistics. To account for uncertainty, a widely used approach consists in constructing a confidence set, known to contain the true mean with high probability, rather than relying solely on a point estimate. When the data are observed sequentially, one might want to update this set as new data-points become available. However, such procedure can compromise the validity of the statistical guarantee, if this was designed for a fixed sample size. To address this issue, Darling and Robbins (1967) introduced the concept of confidence sequence, a data-adaptive sequence of confidence sets whose intersection contains the desired mean with high probability.
Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023) have recently suggested algorithmic approaches that yield some of the tightest confidence sequences for the mean of a bounded real random variable. Both papers propose setting up a series of sequential coin-betting games, one per each mean candidate value $\mu$, where a player sequentially bets on the difference between $\mu$ and the upcoming observation. If $\mu$ matches the true mean, the game is fair, making substantial gains unlikely. A confidence sequence is then obtained by excluding those values $\mu$ that allowed the player to accumulate significant wealth.
This coin-betting approach to mean estimation is a particular instance of a broader algorithmic framework for constructing confidence sequences through sequential hypothesis testing, which can be framed in terms of betting games where at each round the player has to select an e-variable (Shafer, 2021; Ramdas et al., 2022, 2023). E-variables, namely non-negative random variables whose expectation is bounded by one under the tested hypothesis (Grünwald et al., 2024), have recently emerged as a powerful and increasingly popular tool for anytime-valid hypothesis testing. By serving as building blocks for constructing non-negative super-martingales, which can be seen as representing the wealth of a player in a betting game, e-variables naturally lend themselves to game-theoretic interpretations (Shafer and Vovk, 2019; Ramdas and Wang, 2024).
The purpose of this letter is to illustrate that, when estimating the mean of a bounded real random variable, no instantiation of an e-variable-based algorithmic framework can yield strictly better results than the coin-betting approach, in terms of tightness of the resulting confidence sequence. In a sense to be clarified later, coin-betting is optimal, as it represents the “simplest” formulation among those that cannot be strictly outperformed by any other such algorithmic approach. The main novelty of this work is the introduction of a notion of optimality at the level of sets of e-variables. This perspective differs from the existing literature, which is primarily focused on the optimality of an individual e-variable (e.g., log-optimality in Koolen and Grünwald 2022; Grünwald et al. 2024; Larsson et al. 2024) or of a single wealth super-martingale (e.g., admissibility in Ramdas et al. 2022). The author hopes that this new viewpoint will inspire further research.
Notation
Endow the interval $[0,1]$ with the standard topology, and let $\mathcal{P}$ be the set of all Borel probability measures on $[0,1]$. For $x$ in $[0,1]$, $\delta_{x}$ denotes the Dirac unit mass on $x$. For $P\in\mathcal{P}$ and a Borel measurable function $f$ on $[0,1]$, $\langle P,f(X)\rangle$ (or more compactly $\langle P,f\rangle$) denotes the expectation of $f$ under $X\sim P$. Sequences are denoted as $(s_{t})_{t\geq T_{0}}$, with $t$ an integer index and $T_{0}$ its smallest value (typically 00 or $1$). For high probability statements, $\mathbb{P}$ expresses probability with respect to all the randomness involved. For instance, if $(X_{t})_{t\geq 1}$ is a sequence of i.i.d. draws from $P\in\mathcal{P}$, we write $\mathbb{P}(X_{t}\geq 1/2\,,\,\forall t\geq 1)$, with obvious meaning.
2 Algorithmic mean estimation
Let $(X_{t})_{t\geq 1}$ be a sequence of independent random variables drawn from an unknown fixed distribution $P^{\star}\in\mathcal{P}$, whose mean $\mu^{\star}\in(0,1)$ has to be estimated. Let $\mathcal{F}=(\mathcal{F}_{t})_{t\geq 0}$ represent the natural filtration generated by this sequence, where $\mathcal{F}_{t}=\sigma(X_{1},\dots,X_{t})$ captures all information available up to time $t$. Fix a confidence level parameter $\delta\in(0,1)$. A confidence sequence $(S_{t})_{t\geq 1}$ is a sequence of random sets such that the sequence of events ($(\{\mu^{\star}\in S_{t}\})_{t\geq 1}$ is adapted with respect to the filtration $\mathcal{F}$ (i.e., for all $t\geq 1$, $\{\mu^{\star}\in S_{t}\}$ is $\mathcal{F}_{t}$-measurable) and satisfies the property
$\mathbb{P}\big{(}\mu^{\star}\in S_{t}\,,\;\forall t\geq 1\big{)}\geq 1-\delta\,.$
Intuitively, this means that $(S_{t})_{t\geq 1}$ provides a set of plausible values for $\mu^{\star}$ at each time step $t$, while ensuring that the true mean remains in these sets indefinitely with high probability.
2.1 Coin-betting confidence sequence
Recently, Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023) obtained some of the tightest known confidence sequences for bounded random variables via an algorithmic approach. The idea behind both papers is to set up a sequential coin-betting game for each candidate value $\mu\in(0,1)$ for the mean, and to exclude from the confidence sets those values $\mu$ that allow the player to earn unlikely high rewards.
Definition 1 (Coin-betting game)
Fix $\mu\in(0,1)$ and let $I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$. Consider the following sequential procedure. At each round $t\geq 1$, a player
  • 1.
    picks [^1] We implicitly assume a measurable selection of $\beta_{t}$, namely the mapping $(x_{1},\dots x_{t-1})\mapsto\beta_{t}$ is Borel measurable. $\beta_{t}\in I_{\mu}$, based solely on the past observations $x_{1},\dots,x_{t-1}$;
  • 2.
    observes a new data-point $x_{t}\in[0,1]$;
  • 3.
    gets a reward $\log\big{(}1+\beta_{t}(x_{t}-\mu)\big{)}$.
Now, for each $\mu\in(0,1)$, consider a coin-betting game where at each round the player observes a data-point from the sequence $(X_{t})_{t\geq 1}$ drawn from $P^{\star}$. Let $R_{n}(\mu)=\sum_{t=1}^{n}\log(1+\beta_{t}(X_{t}-\mu))$ denote the cumulative reward at round $n$. Then, via standard martingale concentration arguments (which will become clear later) one can show that the sets $S_{n}=\big{\{}\mu\in(0,1)\,:\,R_{n}(\mu)\leq\log\tfrac{1}{\delta}\big{\}}$ define a valid confidence sequence for $\mu^{\star}$.
2.2 Sequential testing game
The coin-betting approach is a particular instance of constructing confidence sequences via a sequential hypothesis testing procedure based on e-variables. We call hypothesis an arbitrary subset $\mathcal{H}$ of $\mathcal{P}$. An e-variable (with respect to $\mathcal{H}$) is any non-negative Borel measurable function $E:[0,1]\to[0,+\infty)$, such that
$\langle P,E\rangle\leq 1\,,\qquad\forall P\in\mathcal{H}\,.$
We denote as $\mathcal{E}_{\mathcal{H}}$ the set of all the e-variables with respect to $\mathcal{H}$, and we call e-class any subset of $\mathcal{E}_{\mathcal{H}}$. We remark that $\mathcal{E}_{\mathcal{H}}$ is never empty, since the constant function one is always an e-variable, independently of $\mathcal{H}$.
Definition 2 (Testing game)
Fix a hypothesis $\mathcal{H}\subseteq\mathcal{P}$ and an e-class $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$. We call $\mathcal{E}$-restricted testing game (on $\mathcal{H}$) the following sequential procedure. At each round $t\geq 1$, a player
  • 1.
    picks [^2] Again, we assume that $E_{t}$ is picked in a measurable fashion, namely the mapping $(x_{1},\dots x_{t})\mapsto E_{t}(x_{t})$ is Borel measurable. an e-variable $E_{t}\in\mathcal{E}$ based solely on the past observations $x_{1},\dots,x_{t-1}$;
  • 2.
    observes a new data-point $x_{t}\in[0,1]$;
  • 3.
    gets a reward $\log E_{t}(x_{t})$.
If $\mathcal{E}=\mathcal{E}_{\mathcal{H}}$, we speak of unrestricted testing game.
The above game is at the core of e-variable testing (see, e.g., Ramdas and Wang 2024 for a recent overview), where one designs a test that rejects the hypothesis $\mathcal{H}$ whenever the total reward earned by the player gets excessively high. This procedure is justified by the fact that, if the data-points observed during the game were independently drawn from $P\in\mathcal{H}$, then the cumulative reward would be unlikely to grow very large.
Lemma 1
Let $\mathcal{H}\subseteq\mathcal{P}$ and consider a sequence $(X_{t}^{\prime})_{t\geq 1}$ of independent draws from some $P\in\mathcal{H}$. Fix $\delta\in(0,1)$ and $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$. Consider an $\mathcal{E}$-restricted testing game, where the observations are the sequence $(X_{t}^{\prime})_{t\geq 1}$. Let $R_{n}=\sum_{t=1}^{n}\log E_{t}(X_{t}^{\prime})$ represent the player’s cumulative reward at round $n$. We have
$\mathbb{P}\big{(}R_{n}\leq\log\tfrac{1}{\delta}\,,\;\forall n\geq 1\big{)}\geq 1-\delta\,.$
Proof 1
The result follows directly from Ville’s inequality, since $M_{n}=\prod_{t=1}^{n}E_{t}(X_{t}^{\prime})$ defines a non-negative super-martingale with respect to the natural filtration of the process $(X_{t}^{\prime})_{t\geq 1}$, with $M_{0}\equiv 1$$\qed$
2.3 Mean estimation via e-variables
We can leverage the testing game of Definition 2 to obtain a confidence sequence for the mean $\mu^{\star}\in(0,1)$ of $P^{\star}$. For $\mu\in(0,1)$, define the hypothesis
$\mathcal{H}_{\mu}=\{P\in\mathcal{P}\,:\,\langle P,X\rangle=\mu\}\,,$ (1)
which contains all probability measures on $[0,1]$ with mean $\mu$. For each $\mu$, we fix an e-class $\mathcal{E}_{\mu}\subseteq\mathcal{E}_{\mathcal{H}_{\mu}}$ and we consider an $\mathcal{E}_{\mu}$-restricted testing game on $\mathcal{H}_{\mu}$, where the player observes $(X_{t})_{t\geq 1}$, the sequence of draws from $P^{\star}$. For each one of these games, denote as $R_{n}(\mu)$ the player’s cumulative reward at round $n$.
Lemma 2
The sequence $(S_{n})_{n\geq 1}$, defined as
$S_{n}=\big{\{}\mu\in(0,1)\;:\;R_{n}(\mu)\leq\log\tfrac{1}{\delta}\big{\}}\,,$
is a confidence sequence for the mean $\mu^{\star}$ of the data-generating probability measure $P^{\star}$.
Proof 2
For each $n\geq 1$, we have that $\{\mu^{\star}\in S_{n}\}=\{R_{n}(\mu^{\star})\leq\log\frac{1}{\delta}\}$, which is a $\mathcal{F}_{n}$-measurable event. Moreover,
$\mathbb{P}\big{(}\mu^{\star}\in S_{n}\,,\;\forall n\geq 1\big{)}=\mathbb{P}\big{(}R_{n}(\mu^{\star})\leq\log\tfrac{1}{\delta}\,,\;\forall n\geq 1\big{)}\geq 1-\delta$
by Lemma 1, as the observations are drawn from $P^{\star}\in\mathcal{H}_{\mu^{\star}}$. $\qed$
We remark that the approach to derive confidence sequences via the coin-betting game, proposed by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023), is a specific instance of the framework that we have just described. Indeed, letting $E_{t}:x\mapsto 1+\beta_{t}(x-\mu)$, it is straightforward to verify that $E_{t}$ is non-negative on $[0,1]$, due to the restriction $\beta_{t}\in I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$ in Definition 1. Moreover, for any $P\in\mathcal{H}_{\mu}$, we have $\langle P,E_{t}\rangle=1$, which implies that $E_{t}$ is an e-variable for $\mathcal{H}_{\mu}$. Finally, the reward in the coin-betting game is precisely equal to $\log E_{t}(x_{t})$. Hence, for the hypothesis $\mathcal{H}_{\mu}$, the coin-betting game matches exactly the testing game of Definition 2, restricted to the coin-betting e-class
$\mathcal{E}^{\mathrm{cb}}_{\mu}=\big{\{}E_{\beta}:x\mapsto 1+\beta(x-\mu)\,,\;\;\beta\in I_{\mu}\big{\}}\,.$ (2)
Ideally, one aims to construct tight confidence sets. However, achieving this depends on the strategy employed in the testing games. For example, stubbornly playing $E_{t}\equiv 1$ at all rounds in every game would produce valid confidence sets ($S_{n}=(0,1)$, for every $n$) of no practical interest. Indeed, a player’s strategy is most effective when it can rapidly increase the cumulative reward, whenever possible, as this allows to fast exclude redundant points from the confidence sets. In short, the highest the rewards, the tightest the confidence sequence. With this in mind, it is clear that carelessly restricting the player’s choice to a subset $\mathcal{E}_{\mu}$ of $\mathcal{E}_{\mathcal{H}_{\mu}}$ could be highly detrimental, as it might force the player to adopt poor strategies (e.g., consider the trivial case $\mathcal{E}_{\mu}=\{1\}$). This point naturally raises the question: “Does restricting the player to the coin-betting e-class (2) loosen the confidence sequence?”. Interestingly, the answer turns out to be negative, as we will demonstrate throughout the remainder of this letter.
3 Majorising e-classes and optimal e-class
We introduce a partial ordering on the set of real functions on $[0,1]$. Given two functions $f,f^{\prime}:[0,1]\to\mathbb{R}$, we say that $f$ majorises $f^{\prime}$, and write $f\succeq f^{\prime}$, if $f(x)\geq f^{\prime}(x)$ for all $x\in[0,1]$. If $f\succeq f^{\prime}$ and there exists a point $x\in[0,1]$ such that $f(x)>f^{\prime}(x)$, we say that $f$ is a strict majoriser of $f^{\prime}$, and we write $f\succ f^{\prime}$.
Definition 3
$E\in\mathcal{E}_{\mathcal{H}}$ is called a maximal e-variable when there is no $E^{\prime}\in\mathcal{E}_{\mathcal{H}}$ such that $E^{\prime}\succ E$.
Next, we introduce a way to compare different e-classes.
Definition 4
Given two e-classes $\mathcal{E}$ and $\mathcal{E}^{\prime}$, we say that $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$ if, for any $E^{\prime}\in\mathcal{E}$, there is an e-variable $E\in\mathcal{E}$ such that $E\succeq E^{\prime}$. We call majorising e-class any e-class that majorises $\mathcal{E}_{\mathcal{H}}$.
Lemma 3
Every majorising e-class contains all the maximal e-variables.
Proof 3
Let $E\in\mathcal{E}_{\mathcal{H}}$ be maximal and $\mathcal{E}$ a majorising e-class. There must be $E^{\prime}\in\mathcal{E}$ such that $E^{\prime}\succeq E$, but since $E$ is maximal it has to be that $E=E^{\prime}$. Hence, $E\in\mathcal{E}$.
The significance of the notion of majorising e-class for our problem is straightforward: if $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$, then any strategy in an $\mathcal{E}^{\prime}$-restricted game can be matched or outperformed (in terms of rewards) by a corresponding strategy in the $\mathcal{E}$-restricted game, regardless of the sequence of observations. This allows us to compare how the restriction to different e-classes affects the testing game of Definition 2.
Definition 5
If a majorising e-class is contained in every other majorising e-class, it is called optimal.
For any $\mathcal{H}$, a majorising e-class always exists, as $\mathcal{E}_{\mathcal{H}}$ itself is a majorising e-class. However, an optimal e-class may not exist. Next, we state a sufficient and necessary condition for the existence of optimal e-classes.
Lemma 4
An optimal e-class exists if, and only if, the set of all maximal e-variables is a majorising e-class. If an optimal e-class exists, it is unique, it corresponds to the set of all maximal e-variables, and it is the only majorising e-class whose elements are all maximal.
Proof 4
Denote as $\hat{\mathcal{E}}$ the set of all the maximal e-variables. Now, assume that there exists an optimal e-class $\mathcal{E}$. First, we show that all its elements must be maximal. For $E\in\mathcal{E}$, consider any element $E^{\prime}\in\mathcal{E}_{\mathcal{H}}$ such that $E^{\prime}\succeq E$. We can construct an e-class $\mathcal{E}^{\prime}$ replacing $E$ with $E^{\prime}$ in $\mathcal{E}$, namely =($\mathcal{E}^{\prime}=(\mathcal{E}\setminus\{E\})\cup\{E^{\prime}\}$. Since $E^{\prime}\succeq E$, it is clear that $\mathcal{E}^{\prime}$ majorises $\mathcal{E}$. It follows that $\mathcal{E}^{\prime}$ is a majorising e-class, as $\mathcal{E}$ is. Since $\mathcal{E}$ is optimal, $\mathcal{E}\subseteq\mathcal{E}^{\prime}$, which implies $E=E^{\prime}$. In particular, $E$ does not have any strict majoriser, and so it is maximal. Hence, $\mathcal{E}\subseteq\hat{\mathcal{E}}$. As $\mathcal{E}\supseteq\hat{\mathcal{E}}$ by Lemma 3, we conclude that $\mathcal{E}=\hat{\mathcal{E}}$, and so $\hat{\mathcal{E}}$ is a majorising e-class.
Conversely, assume that $\hat{\mathcal{E}}$ is a majorising e-class. Let $\mathcal{E}$ be any other majorising e-class. By Lemma 3, $\hat{\mathcal{E}}\subseteq\mathcal{E}$. So, $\hat{\mathcal{E}}$ is contained in all the majorising e-classes, and hence it is optimal.
Now, the final statements are a trivial consequence of what shown above and Lemma 3$\qed$
Let us emphasise once more that, from our discussion thus far, it is clear that restricting the testing game of Definition 2 to a majorising e-class does not hinder the performance of the player, as for any unrestricted strategy $(E_{t})_{t\geq 1}$ they can always pick a restricted strategy $(E_{t}^{\prime})_{t\geq 1}$, whose cumulative rewards inevitably match or outperform those of $(E_{t})_{t\geq 1}$, regardless of the sequence of observations. From a practical perspective, identifying the optimal e-class, when it exists, greatly simplifies the design of an effective strategy by narrowing the player’s choice to the best possible e-variables. Specifically, if the optimal e-class exists and a player chooses an e-variable $E_{t}$ outside of it, they could always have picked an alternative $E_{t}^{\prime}\succ E_{t}$, within the optimal e-class, whose reward is never worse than that of $E_{t}$ and is strictly higher for at least one possible value that $x_{t}$ might take. Conversely, when a player selects an e-variable from the optimal e-class, no other choice can be guaranteed to be better before observing $x_{t}$, since the player’s pick is a maximal e-variable. As a straightforward consequence, the optimal approach to obtain confidence sequences consists in restricting the testing game to the optimal e-class, for each mean candidate value $\mu$. We will show next that this coincides precisely with the coin-betting formulation.
4 Optimality of the coin-betting e-class
For any $\mu\in(0,1)$, define the hypothesis $\mathcal{H}_{\mu}$ as in (1). We now show that the optimal e-class for $\mathcal{H}_{\mu}$ exists and coincides with the coin-betting e-class $\mathcal{E}^{\mathrm{cb}}_{\mu}$, defined in (2).
Lemma 5
Fix $\mu\in(0,1)$ and consider the function $F_{\mu}:[0,1]\to[1,+\infty)$ defined as
$F_{\mu}\;:\;x\mapsto\begin{cases}1+\frac{1}{\mu}(x-\mu)&\text{if $x\geq\mu$;}\\ 1+\frac{1}{\mu-1}(x-\mu)&\text{if $x<\mu$.}\end{cases}$
For any $x\in[0,1]$ there is $P_{x}\in\mathcal{H}_{\mu}$ such that $P_{x}(\{x\})=1/F_{\mu}(x)>0$. Moreover, $F_{\mu}\succeq E$ for every $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$.
Proof 5
For $x\geq\mu$, let $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x})\delta_{0}$. Then, $P_{x}\in\mathcal{H}_{\mu}$ and $P_{x}(\{x\})=\mu/x=1/F_{\mu}(x)$. With a similar reasoning, if $x<\mu$ we can find a measure $P_{x}$ in $\mathcal{H}_{\mu}$, supported on $\{x,1\}$, with mass $1/F_{\mu}(x)$ on $x$.
Now, to check that $F$ majorises all the e-variables, fix $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$ and $x\in[0,1]$. Consider a $P_{x}\in\mathcal{H}_{\mu}$ with mass $1/F_{\mu}(x)$ on $x$. Then, $1\geq\langle P_{x},E\rangle\geq P_{x}(x)E(x)=E(x)/F_{\mu}(x)$, and we conclude. $\qed$
Theorem 1
For any $\mu\in(0,1)$, the coin-betting e-class $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is the optimal e-class for $\mathcal{H}_{\mu}$.
Proof 6
First, let us show that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class. Fix an arbitrary $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$. Define the sets
$B_{0}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in[0,\mu]}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\qquad\text{and}\qquad B_{1}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in[\mu,1]}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\,,$
where we recall that $I_{\mu}=[(1-\mu)^{-1},\mu^{-1}]$ and $E_{\beta}:x\mapsto 1+\beta(x-\mu)$. Both sets are closed and convex (as they are intersections of closed and convex sets). By Lemma 5, $E_{(\mu-1)^{-1}}\in B_{0}$ and $E_{\mu^{-1}}\in B_{1}$, so $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu]$, for some $\beta_{0}$ and $\beta_{1}$ in $I_{\mu}$. We will now show that $B_{0}\cap B_{1}\neq\varnothing$, or equivalently that $\beta_{0}\geq\beta_{1}$. Assume that this was not the case and $\beta_{0}<\beta_{1}$. Let $\beta^{\star}\in(\beta_{0},\beta_{1})$. Then, $\beta^{\star}\notin B_{0}$ and $\beta^{\star}\notin B_{1}$. In particular, there are $x_{0}\in[0,\mu]$ and $x_{1}\in[\mu,1]$ such that $E(x_{0})>E_{\beta^{\star}}(x_{0})$ and $E(x_{1})>E_{\beta^{\star}}(x_{1})$. Notice that since $\delta_{\mu}\in\mathcal{H}_{\mu}$, it must be that $E(\mu)\leq 1$, and so $x_{0}<\mu$ and $x_{1}>\mu$. Hence, $\mu\in(x_{0},x_{1})$ and so there is $\hat{P}\in\mathcal{H}_{\mu}$ with support $\{x_{0},x_{1}\}$. Notice that $\langle\hat{P},E_{\beta^{\star}}\rangle=1+\beta^{\star}(\langle\hat{P},X\rangle-\mu)=1$. But $E$ is strictly larger than $E_{\beta^{\star}}$ on $\mathrm{Supp}(\hat{P})$, and so $\langle\hat{P},E\rangle>1$, which is a contradiction since $E$ is an e-variable. We conclude that $\beta_{0}\geq\beta_{1}$ and so there exists $\hat{\beta}\in B_{0}\cap B_{1}$. By construction, $E_{\hat{\beta}}\in\mathcal{E}^{\mathrm{cb}}_{\mu}$ and $E_{\hat{\beta}}\succeq E$. Because the choice of $E$ was arbitrary, $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class.
Now that we have established that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class, by Lemma 4 we only need to show that all its elements are maximal to conclude that it is optimal. Fix $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$, and consider an e-variable $\hat{E}\in\mathcal{E}_{\mathcal{H}_{\mu}}$ such that $\hat{E}\succeq E$. Fix any $x\in[0,1]$. By Lemma 5, there is $P\in\mathcal{H}_{\mu}$ such that $P(\{x\})>0$. Since $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$, we have $\langle P,E\rangle=1$, and so $0\leq P(\{x\})(\hat{E}(x)-E(x))\leq\langle P,\hat{E}-E\rangle=\langle P,\hat{E}\rangle-1\leq 0$. Since $P(\{x\})>0$, we get $\hat{E}(x)=E(x)$ and, $x$ being arbitrary, $\hat{E}=E$. Hence $E$ is maximal, as it has no strict majoriser. $\qed$
5 Perspectives
Theorem 1 gives a rigorous sense to the claim that the coin-betting formulation is optimal among the e-variable-based approaches to build confidence sequences for $\mu^{\star}$. To formalise this assertion, we introduced the notions of majorising and optimal e-classes, which, in the author’s view, are of independent interest in the context of sequential testing, beyond the scope of this letter. The main novelty of these concepts lies in defining “optimality” in terms of e-classes, rather than individual e-variables. This perspective contrasts with the notion of log-optimality of a single e-variable with respect to an alternative hypothesis, widely discussed in the literature (e.g., Koolen and Grünwald, 2022; Grünwald et al., 2024; Larsson et al., 2024).
Rather than log-optimality, what probably best matches our way of “ranking” e-variables is the idea of admissibility introduced by Ramdas et al. (2022). Roughly speaking, let us say that a wealth-process (namely, a non-negative adapted process starting at one that is a super-martingale under every $P\in\mathcal{H}$) is admissible if it is not dominated by any other wealth-process. This concept is tightly related to our notion of maximal e-variables, as it is not hard to see that any wealth-process $(M_{n})_{n\geq 1}$ on $\mathcal{H}_{\mu}$ is admissible if, and only if, it can be associated with a sequence $(E_{t})_{t\geq 1}$ of maximal e-variables such that $M_{n}=\prod_{t=1}^{n}E_{t}(X_{t}^{\prime})$, with $(X_{t}^{\prime})_{t\geq 1}$ a sequence generated by some $P\in\mathcal{H}_{\mu}$. Thus, in our setting, Theorem 1 provides a characterisation of the set of all admissible wealth-processes for $\mathcal{H}_{\mu}$. Let us emphasise that the tools from Ramdas et al. (2022) alone are insufficient to achieve this, as the necessary conditions for admissibility provided by their work rely on the existence of a “reference measure” for the hypothesis, [^3] A hypothesis $\mathcal{H}$ admits a “reference measure” if there exists a Borel measure $Q$ such that $P\ll Q$ for all $P\in\mathcal{H}$. a property that $\mathcal{H}_{\mu}$ does not satisfy. More precisely, their results are enough to prove that every e-variable in the coin-betting e-class is maximal, but not that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class. This point is subtle but essential: establishing the optimality of coin-betting requires studying the properties of a set of e-variables (or equivalently, wealth-processes) rather than focusing on single e-variables or wealth-processes individually. Crucially, it is precisely the fact that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class that ensures that nothing is lost by relying on the coin-betting approach.
To the author’s knowledge, this is the first work to rigorously examine the optimality of the coin-betting formulation, as previous discussion on the topic has relied solely on heuristic arguments. For instance, Waudby-Smith and Ramdas (2023) justify restricting to the coin-betting e-class by noticing that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is precisely the set ¯$\bar{\mathcal{E}}_{\mathcal{H}_{\mu}}$ of e-variables whose expectation equals one under every $P\in\mathcal{H}_{\mu}$. However, this property does not generally imply the optimality of the e-class, as it merely indicates that all its elements are maximal. As a simple counterexample, consider $\mathcal{H}^{\prime}_{\mu}=\{P\in\mathcal{P}\,:\,\langle P,X\rangle\leq\mu\}$. Then, ¯$\bar{\mathcal{E}}_{\mathcal{H}^{\prime}_{\mu}}=\{1\}$, while the optimal e-class exists and consists of all functions in the form $E_{\beta}:x\mapsto 1+\beta(x-\mu)$, with $\beta\in[0,\mu^{-1}]$.
To simplify the exposition, we have focused on random variables valued in $[0,1]$. However, all the results presented can be extended to any bounded closed interval. More broadly, the author anticipates that similar conclusions hold for more general settings, such as estimating the mean of bounded vector-valued random variables. Addressing this case requires more advanced analytic tools and lies beyond the scope of this letter.
As a final remark, this paper does not aim to delve into the design of effective coin-betting strategies for obtaining tight confidence sequences. For this, we refer the interested reader to the thorough analysis and discussion by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023).
Acknowledgements
The author would like to thank Gergely Neu, Peter Daniel Grünwald, Nishant Ajay Mehta, Hamish Edward Flynn, and Claudia Maria Chanu, for the insightful discussions that inspired this work. The author acknowledges the use of GPT-4o during the redaction of this letter, in order to polish the presentation. All AI-generated text was subsequently reviewed and edited by the author, who takes full responsibility for the content of this manuscript. This project was funded by the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation programme (grant agreement 950180).