### v2

# On the optimality of coin-betting for mean estimation

Abstract
Confidence sequences are sequences of confidence sets that adapt to incoming data while maintaining validity. Recent advances have introduced an algorithmic formulation for constructing some of the tightest confidence sequences for the mean of bounded real random variables. These approaches use a coin-betting framework, where a player sequentially bets on differences between potential mean values and observed data. This work discusses the optimality of such coin-betting formulation among algorithmic frameworks building on e-variables methods to test and estimate the mean of bounded random variables.
1 Introduction
Estimating the mean of a random variable from empirical observations is a classical problem in statistics. To account for uncertainty, a widely used approach consists in constructing a confidence set, known to contain the true mean with high probability, rather than relying solely on a point estimate. When the data are observed sequentially, one might want to update this set as new data-points become available. However, such procedure can compromise the validity of the statistical guarantee, if this was designed for a fixed sample size. To address this issue, Darling and Robbins (1967) introduced the concept of confidence sequence, a data-adaptive sequence of confidence sets whose intersection contains the desired mean with high probability.
Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023) have recently suggested algorithmic approaches that yield some of the tightest confidence sequences for the mean of a bounded real random variable. Both papers propose setting up a series of sequential coin-betting games, one per each mean candidate value $\mu$, where a player sequentially bets on the difference between $\mu$ and the upcoming observation. If $\mu$ matches the true mean, the game is fair, making substantial gains unlikely. A confidence sequence is then obtained by excluding those values $\mu$ that allowed the player to accumulate significant wealth.
This coin-betting approach to mean estimation is a particular instance of a broader algorithmic framework for constructing confidence sequences through sequential hypothesis testing, which can be framed in terms of betting games where at each round the player has to select an e-variable (Shafer, 2021; Ramdas et al., 2022a, 2023). E-variables, non-negative random variables whose expectation is bounded by one under the tested hypothesis (Grünwald et al., 2024), have recently emerged as a powerful and increasingly popular tool for anytime-valid hypothesis testing. By serving as building blocks for constructing non-negative super-martingales, which can be seen as representing the wealth of a player in a betting game, e-variables naturally lend themselves to game-theoretic interpretations (Shafer and Vovk, 2019; Ramdas and Wang, 2024).
The first part of this work focuses on round-wise testing-by-betting, a scenario where the player iteratively picks a single-round e-variable, whose choice can depend on the past observations. Our main goal is to illustrate that, when estimating the mean of a bounded real random variable, no instantiation of such round-wise testing-by-betting game can yield strictly better results than the coin-betting approach, in terms of tightness of the resulting confidence sequence. In a sense to be clarified later, coin-betting is optimal, as it represents the “simplest” formulation among those that cannot be strictly outperformed by any other such betting approach. One main novelty of this work is the introduction of a notion of optimality at the level of sets of e-variables. This perspective differs from much of the existing literature, which is primarily focused on the optimality of an individual e-variable or wealth process (e.g., log-optimality in Koolen and Grünwald 2022; Grünwald et al. 2024; Larsson et al. 2024, or admissibility in Ramdas et al. 2022a).
An alternative way to frame the first part of this work is as a characterisation of the family of single-round e-variables for testing the mean of a bounded random variable. In a similar spirit, since the appearance of the first version of this manuscript, two works have derived more general results for testing hypotheses defined by linear constraints (Clerico, 2024; Larsson et al., 2025), of which the present setting is a particular case. However, the approach we present here is direct and tailored to the simple setting considered, making it valuable for building intuition, while the more general results are significantly more abstract.
We remark that testing via sequential betting with single-round e-variables is not the most general form of sequential testing with e-variables, which typically relies on multi-round e-variables and e-processes (Shafer, 2021; Koolen and Grünwald, 2022; Ramdas et al., 2022b; Ramdas and Wang, 2024). These allow for testing hypotheses over the entire data sequence, including assumptions about the dependence structure (e.g., i.i.d. or fixed conditional mean). In such cases, restricting the player to select a single-round e-variable at each step can be limiting (Koolen and Grünwald, 2022; Ramdas et al., 2022b). The second part of this work considers this broader setting. For technical convenience, we only focus on variables with finite discrete support, though we conjecture our conclusions hold more generally. We establish that when the sequence has fixed conditional mean, coin betting remains optimal even among testing methods based on multi-round e-variables and e-processes. However, we also show that this optimality result no longer holds under the more restrictive hypothesis of i.i.d. observations.
Notation
We endow any Borel set $\mathcal{Z}\subseteq\mathbb{R}^{d}$ with the standard topology, and we denote as $\mathcal{P}_{\mathcal{Z}}$ the set of Borel probability measures on $\mathcal{Z}$. For $z\in\mathcal{Z}$, $\delta_{z}$ denotes the Dirac unit mass on $z$. For $P\in\mathcal{P}_{\mathcal{Z}}$ and a Borel measurable function $f$ on $\mathcal{Z}$, $\mathbb{E}_{P}[f(Z)]$ (or more compactly $\mathbb{E}_{P}[f]$) denotes the expectation of $f$ under $Z\sim P$. Given a sub-sigma-field $\mathcal{G}$, $\mathbb{E}_{P}[Z|\mathcal{G}]$ denotes the conditional expectation. Typically, $\mathcal{G}$ will be the sigma-field generated by some random variable $X$. In such case, we will write $\mathbb{E}_{P}[Z|X]$.
We will also consider measures on product spaces. For $T\geq 1$, we write $\mathcal{P}_{\mathcal{Z}^{T}}$ the set of Borel probability measure on $\mathcal{Z}^{T}$ (endowed with the product topology). We will used $\otimes$ to denote the direct product of measures. For instance, given $P$ and $Q$ in $\mathcal{P}_{\mathcal{Z}}$, $P\otimes Q$ will be the element of $\mathcal{P}_{Z^{2}}$ that encode the law of $(Z_{1},Z_{2})$, where $Z_{1}\sim P$ and $Z_{2}\sim Q$ are independent.
For any two given integers $s$ and $t$, with $s\leq t$, $[s:t]$ denotes the set of integers between $s$ and $t$ (included). For an integer $T$, given a vector $(x_{1},\dots,x_{T})$, we often represent it compactly as $x^{T}$ (upper indices). At times, we will also use the notation $x^{t:T}$ (with $t\in[1:T]$), to denote the vector $(x_{t},x_{t+1},\dots,x_{T})$. Sequences are denoted as $(s_{t})_{t\geq T_{0}}$, with $t$ an integer index and $T_{0}$ its smallest value (typically 00 or $1$). Sometimes we will also use the notation $s^{T_{0}:\infty}$ to denote $(s_{t})_{t\geq T_{0}}$, or simply $s^{\infty}$ if $T_{0}$ is clear from the context. For high probability statements, $\mathbb{P}$ expresses probability with respect to all the randomness involved. For instance, if $(X_{t})_{t\geq 1}$ is a sequence of i.i.d. draws from $P\in\mathcal{P}_{\mathcal{X}}$, we write $\mathbb{P}(X_{t}\geq 1/2\,,\,\forall t\geq 1)$, with obvious meaning.
2 Algorithmic mean estimation
Fix a Borel set $\mathcal{X}\subseteq\mathbb{R}$. Let $(X_{t})_{t\geq 1}$ be a sequence of independent random variables drawn from an unknown fixed distribution $P^{\star}\in\mathcal{P}_{\mathcal{X}}$, whose mean $\mu^{\star}\in(0,1)$ has to be estimated. Let $\mathcal{F}=(\mathcal{F}_{t})_{t\geq 0}$ represent the natural filtration generated by this sequence, where $\mathcal{F}_{t}=\sigma(X_{1},\dots,X_{t})$ captures all information available up to time $t$. Fix a confidence level parameter $\delta\in(0,1)$. A confidence sequence $(S_{t})_{t\geq 1}$ is a sequence of random sets such that the sequence of events ($(\{\mu^{\star}\in S_{t}\})_{t\geq 1}$ is adapted with respect to the filtration $\mathcal{F}$ (i.e., for all $t\geq 1$, $\{\mu^{\star}\in S_{t}\}$ is $\mathcal{F}_{t}$-measurable) and satisfies the property
$\mathbb{P}\big{(}\mu^{\star}\in S_{t}\,,\;\forall t\geq 1\big{)}\geq 1-\delta\,.$
Intuitively, this means that $(S_{t})_{t\geq 1}$ provides a set of plausible values for $\mu^{\star}$ at each time step $t$, while ensuring that the true mean remains in these sets indefinitely with high probability.
We stress here that, for the time being, we assume that the sequence $(X_{t})_{t\geq 1}$ is i.i.d., and we shall not challenge this assumption, regardless of whether or not the observed data appear to exhibit such behaviour. In a way, this perspective aligns naturally with the goal of constructing confidence sequences, as simultaneously testing for the mean and the i.i.d. assumption might lead to rejecting everything and returning empty confidence sets because “the data do not look i.i.d.”. Under the i.i.d. model, mean estimation is well posed: we assume the existence of a fixed distribution $P^{\star}$ independently generating each observation, and our task is to estimate its mean. We remark that the independence assumption could be relaxed by only asking that each $X_{t}$ has fixed conditional mean $\mu^{\star}$ (to be estimated) given the past $\mathcal{F}_{t-1}$. The arguments we present next for the i.i.d. setting carry over with essentially no change to this less restrictive case. However, to keep the exposition clearer, we focus here solely on the independent case. The conditional setting will be addressed in the second part of this work, where we consider the more general scenario in which dependencies between observations are not assumed but are instead challenged by the statistical test.
2.1 Coin-betting confidence sequence
Recently, Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023) obtained some of the tightest known confidence sequences for bounded random variables via an algorithmic approach. The idea behind both papers is to set up a sequential coin-betting game for each candidate value $\mu\in(0,1)$ for the mean, and to exclude from the confidence sets those values $\mu$ that allow the player to earn unlikely high rewards.
Definition 1 (Coin-betting game).
Fix $\mu\in(0,1)$ and let $I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$. Consider the following sequential procedure. At each round $t\geq 1$, a player
  • picks [^2] We implicitly assume a measurable selection of $\beta_{t}$, namely the mapping $(x_{1},\dots x_{t-1})\mapsto\beta_{t}$ is Borel measurable. $\beta_{t}\in I_{\mu}$, based solely on the past observations $x_{1},\dots,x_{t-1}$;
  • observes a new data-point $x_{t}\in[0,1]$;
  • gets a reward $\log\big{(}1+\beta_{t}(x_{t}-\mu)\big{)}$.
Now, for each $\mu\in(0,1)$, consider a coin-betting game where at each round the player observes a data-point from the sequence $(X_{t})_{t\geq 1}$ drawn from $P^{\star}$. Let $R_{n}(\mu)=\sum_{t=1}^{n}\log(1+\beta_{t}(X_{t}-\mu))$ denote the cumulative reward at round $n$. Then, via standard martingale concentration arguments (which will become clear later) one can show that the sets $S_{n}=\big{\{}\mu\in(0,1)\,:\,R_{n}(\mu)\leq\log\tfrac{1}{\delta}\big{\}}$ define a valid confidence sequence for $\mu^{\star}$.
2.2 Sequential testing game
The coin-betting approach is a particular instance of constructing confidence sequences via a sequential hypothesis testing procedure based on e-variables. Here, we call hypothesis an arbitrary subset $\mathcal{H}$ of $\mathcal{P}_{[0,1]}$. An e-variable (for $\mathcal{H}$) is any non-negative Borel measurable function $E:[0,1]\to[0,+\infty)$, such that
$\mathbb{E}_{P}[E]\leq 1\,,\qquad\forall P\in\mathcal{H}\,.$
We denote as $\mathcal{E}_{\mathcal{H}}$ the set of all the e-variables with respect to $\mathcal{H}$, and we call e-class any subset of $\mathcal{E}_{\mathcal{H}}$. We remark that $\mathcal{E}_{\mathcal{H}}$ is never empty, since the constant function one is always an e-variable, independently of $\mathcal{H}$.
Definition 2 (Testing-by-betting game).
Fix a hypothesis $\mathcal{H}\subseteq\mathcal{P}_{[0,1]}$ and a non-empty e-class $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$. We call $\mathcal{E}$-restricted testing-by-betting game (on $\mathcal{H}$) the following sequential procedure. At each round $t\geq 1$,
  • the player picks [^3] Again, we assume that $E_{t}$ is picked in a measurable fashion, namely the mapping $(x_{1},\dots x_{t})\mapsto E_{t}(x_{t})$ is Borel measurable. an e-variable $E_{t}\in\mathcal{E}$ based solely on the past observations $x_{1},\dots,x_{t-1}$;
  • the player observes a new data-point $x_{t}\in[0,1]$;
  • the player earns a reward $\log E_{t}(x_{t})$.
If $\mathcal{E}=\mathcal{E}_{\mathcal{H}}$, we speak of unrestricted testing-by-betting game.
The above game is an instance of e-variable testing (see Ramdas and Wang, 2024, for a recent overview), where one designs a test that rejects the hypothesis $\mathcal{H}$ whenever the total reward earned by the player gets excessively high. This procedure is justified by the fact that, if the data-points observed during the game were independently drawn from $P\in\mathcal{H}$, then the cumulative reward would be unlikely to grow very large.
Lemma 1.
Let $\mathcal{H}\subseteq\mathcal{P}_{[0,1]}$ and consider a sequence $(X_{t}^{\prime})_{t\geq 1}$ of independent draws from some $P\in\mathcal{H}$. Fix $\delta\in(0,1)$ and $\mathcal{E}\subseteq\mathcal{E}_{\mathcal{H}}$. Consider an $\mathcal{E}$-restricted testing-by-betting game, where the observations are the sequence $(X_{t}^{\prime})_{t\geq 1}$. Let $R_{n}=\sum_{t=1}^{n}\log E_{t}(X_{t}^{\prime})$ represent the player’s cumulative reward at round $n$. We have
$\mathbb{P}\big{(}R_{n}\leq\log\tfrac{1}{\delta}\,,\;\forall n\geq 1\big{)}\geq 1-\delta\,.$
Proof.
The result follows directly from Ville’s inequality, since $M_{n}=\prod_{t=1}^{n}E_{t}(X_{t}^{\prime})$ defines a non-negative super-martingale with respect to the natural filtration of the process $(X_{t}^{\prime})_{t\geq 1}$, with $M_{0}\equiv 1$. ∎
2.3 Mean estimation via sequential betting
We can leverage the testing-by-betting game of Definition 2 to obtain a confidence sequence for the mean $\mu^{\star}\in(0,1)$ of $P^{\star}$. For $\mu\in(0,1)$, define the hypothesis
$\mathcal{H}_{\mu}=\{P\in\mathcal{P}_{[0,1]}\,:\,\mathbb{E}_{P}[X]=\mu\}\,,$ (1)
which contains all probability measures on $[0,1]$ with mean $\mu$. For each $\mu$, we fix an e-class $\mathcal{E}_{\mu}\subseteq\mathcal{E}_{\mathcal{H}_{\mu}}$ and we consider an $\mathcal{E}_{\mu}$-restricted testing-by-betting game on $\mathcal{H}_{\mu}$, where the player observes $(X_{t})_{t\geq 1}$, the sequence of draws from $P^{\star}$. For each one of these games, denote as $R_{n}(\mu)$ the player’s cumulative reward at round $n$.
Proposition 1.
The sequence $(S_{n})_{n\geq 1}$, defined as
$S_{n}=\big{\{}\mu\in(0,1)\;:\;R_{n}(\mu)\leq\log\tfrac{1}{\delta}\big{\}}\,,$
is a confidence sequence for the mean $\mu^{\star}$ of the data-generating probability measure $P^{\star}$.
Proof.
For each $n\geq 1$, we have that $\{\mu^{\star}\in S_{n}\}=\{R_{n}(\mu^{\star})\leq\log\frac{1}{\delta}\}$, which is a $\mathcal{F}_{n}$-measurable event. Moreover,
$\mathbb{P}\big{(}\mu^{\star}\in S_{n}\,,\;\forall n\geq 1\big{)}=\mathbb{P}\big{(}R_{n}(\mu^{\star})\leq\log\tfrac{1}{\delta}\,,\;\forall n\geq 1\big{)}\geq 1-\delta$
by Lemma 1, as the observations are drawn from $P^{\star}\in\mathcal{H}_{\mu^{\star}}$. ∎
We remark that the approach to derive confidence sequences via the coin-betting game, proposed by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023), is a specific instance of the framework that we have just described. Indeed, letting $E_{t}:x\mapsto 1+\beta_{t}(x-\mu)$, it is straightforward to verify that $E_{t}$ is non-negative on $[0,1]$, due to the restriction $\beta_{t}\in I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$ in Definition 1. Moreover, for any $P\in\mathcal{H}_{\mu}$, we have $\mathbb{E}_{P}[E_{t}]=1$, which implies that $E_{t}$ is an e-variable for $\mathcal{H}_{\mu}$. Finally, the reward in the coin-betting game is precisely equal to $\log E_{t}(x_{t})$. Hence, for the hypothesis $\mathcal{H}_{\mu}$, the coin-betting game matches exactly the testing game of Definition 2, restricted to the coin-betting e-class
$\mathcal{E}^{\mathrm{cb}}_{\mu}=\big{\{}E_{\beta}:x\mapsto 1+\beta(x-\mu)\,,\;\;\beta\in I_{\mu}\big{\}}\,.$ (2)
Ideally, one aims to construct tight confidence sets. However, achieving this depends on the strategy employed in the testing games. For example, stubbornly playing $E_{t}\equiv 1$ at all rounds in every game would produce valid confidence sets ($S_{n}=(0,1)$, for every $n$) of no practical interest. Indeed, a player’s strategy is most effective when it can rapidly increase the cumulative reward, whenever possible, as this allows to fast exclude redundant points from the confidence sets. In short, the highest the rewards, the tightest the confidence sequence. With this in mind, it is clear that carelessly restricting the player’s choice to a subset $\mathcal{E}_{\mu}$ of $\mathcal{E}_{\mathcal{H}_{\mu}}$ could be highly detrimental, as it might force the player to adopt poor strategies (e.g., consider the trivial case $\mathcal{E}_{\mu}=\{1\}$). This point naturally raises the question: “Does restricting the player to the coin-betting e-class (2) loosen the confidence sequence?”. Interestingly, for the round-wise testing-by-betting framework that we are considering the answer turns out to be negative.
3 Majorising e-classes and optimal e-class
We introduce a partial ordering on the set of real functions on a set $\mathcal{X}$. Given two functions $f,f^{\prime}:\mathcal{X}\to\mathbb{R}$, we say that $f$ majorises $f^{\prime}$, and write $f\succeq f^{\prime}$, if $f(x)\geq f^{\prime}(x)$ for all $x\in\mathcal{X}$. If $f\succeq f^{\prime}$ and there exists a point $x\in\mathcal{X}$ such that $f(x)>f^{\prime}(x)$, we say that $f$ is a strict majoriser of $f^{\prime}$, and we write $f\succ f^{\prime}$.
Definition 3.
$E\in\mathcal{E}_{\mathcal{H}}$ is called a maximal e-variable when there is no $E^{\prime}\in\mathcal{E}_{\mathcal{H}}$ such that $E^{\prime}\succ E$.
Next, we introduce a way to compare different e-classes.
Definition 4.
Given two e-classes $\mathcal{E}$ and $\mathcal{E}^{\prime}$, we say that $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$ if, for any $E^{\prime}\in\mathcal{E}^{\prime}$, there is an e-variable $E\in\mathcal{E}$ such that $E\succeq E^{\prime}$. We call majorising e-class any e-class that majorises $\mathcal{E}_{\mathcal{H}}$.
Lemma 2.
Every majorising e-class contains all the maximal e-variables.
Proof.
Let $E\in\mathcal{E}_{\mathcal{H}}$ be maximal and $\mathcal{E}$ a majorising e-class. There must be $E^{\prime}\in\mathcal{E}$ such that $E^{\prime}\succeq E$, but since $E$ is maximal it has to be that $E=E^{\prime}$. Hence, $E\in\mathcal{E}$. ∎
The significance of the notion of majorising e-class for our problem is straightforward: if $\mathcal{E}$ majorises $\mathcal{E}^{\prime}$, then any strategy in an $\mathcal{E}^{\prime}$-restricted game can be matched or outperformed (in terms of rewards) by a corresponding strategy in the $\mathcal{E}$-restricted game, regardless of the sequence of observations. This allows us to compare how the restriction to different e-classes affects the testing-by-betting game of Definition 2.
Definition 5.
If a majorising e-class is contained in every other majorising e-class, we call it optimal.
For any $\mathcal{H}$, a majorising e-class always exists, as $\mathcal{E}_{\mathcal{H}}$ itself is a majorising e-class. However, an optimal e-class may not exist. [^4] We refer the reader to the follow-up work Clerico (2024) for an example of non-existence of the optimal e-class. Proposition 2 therein shows that =$\mathcal{H}=\{P\in\mathcal{P}_{[0,1]}\,:\,P(\{0\})\geq 1/2\}\cup\{U_{[0,1]}\}$ (with $U_{[0,1]}$ the uniform distribution on $[0,1]$) is a hypothesis for which the set of maximal e-variables is not a majorising e-class. Hence, $\mathcal{H}$ does not admit an optimal e-class. Next, we state a sufficient and necessary condition for the existence of optimal e-classes.
Lemma 3.
An optimal e-class exists if, and only if, the set of all maximal e-variables is a majorising e-class. If an optimal e-class exists, it is unique, it corresponds to the set of all maximal e-variables, and it is the only majorising e-class whose elements are all maximal.
Proof.
Denote as $\hat{\mathcal{E}}$ the set of all the maximal e-variables. Now, assume that there exists an optimal e-class $\mathcal{E}$. First, we show that all its elements must be maximal. For $E\in\mathcal{E}$, consider any element $E^{\prime}\in\mathcal{E}_{\mathcal{H}}$ such that $E^{\prime}\succeq E$. We can construct an e-class $\mathcal{E}^{\prime}$ replacing $E$ with $E^{\prime}$ in $\mathcal{E}$, namely =($\mathcal{E}^{\prime}=(\mathcal{E}\setminus\{E\})\cup\{E^{\prime}\}$. Since $E^{\prime}\succeq E$, it is clear that $\mathcal{E}^{\prime}$ majorises $\mathcal{E}$. It follows that $\mathcal{E}^{\prime}$ is a majorising e-class, as $\mathcal{E}$ is. Since $\mathcal{E}$ is optimal, $\mathcal{E}\subseteq\mathcal{E}^{\prime}$, which implies $E=E^{\prime}$. In particular, $E$ does not have any strict majoriser, and so it is maximal. Hence, $\mathcal{E}\subseteq\hat{\mathcal{E}}$. As $\mathcal{E}\supseteq\hat{\mathcal{E}}$ by Lemma 2, we conclude that $\mathcal{E}=\hat{\mathcal{E}}$, and so $\hat{\mathcal{E}}$ is a majorising e-class.
Conversely, assume that $\hat{\mathcal{E}}$ is a majorising e-class. Let $\mathcal{E}$ be any other majorising e-class. By Lemma 2, $\hat{\mathcal{E}}\subseteq\mathcal{E}$. So, $\hat{\mathcal{E}}$ is contained in all the majorising e-classes, and hence it is optimal.
Now, the final statements are a trivial consequence of what was shown above and Lemma 2. ∎
Let us emphasise once more that, from our discussion thus far, it is clear that restricting the testing-by-betting game of Definition 2 to a majorising e-class does not hinder the performance of the player, as for any unrestricted strategy $(E_{t})_{t\geq 1}$ they can always pick a restricted strategy $(E_{t}^{\prime})_{t\geq 1}$, whose cumulative rewards inevitably match or outperform those of $(E_{t})_{t\geq 1}$, regardless of the sequence of observations. From a practical perspective, identifying the optimal e-class, when it exists, greatly simplifies the design of an effective strategy by narrowing the player’s choice to the best possible e-variables. Specifically, if the optimal e-class exists and a player chooses an e-variable $E_{t}$ outside of it, they could always have picked an alternative $E_{t}^{\prime}\succ E_{t}$, within the optimal e-class, whose reward is never worse than that of $E_{t}$ and is strictly higher for at least one possible value that $x_{t}$ might take. Conversely, when a player selects an e-variable from the optimal e-class, no other choice can be guaranteed to be better before observing $x_{t}$, since the player’s pick is a maximal e-variable. As a straightforward consequence, within this round-wise testing-by-betting approach, the optimal approach to obtain confidence sequences consists in restricting game to the optimal e-class, for each mean candidate value $\mu$. We show next that this coincides precisely with the coin-betting formulation.
4 Optimality of the coin-betting e-class
For any $\mu\in(0,1)$, define the hypothesis $\mathcal{H}_{\mu}$ as in (1). We now show that the optimal e-class for $\mathcal{H}_{\mu}$ exists and coincides with the coin-betting e-class $\mathcal{E}^{\mathrm{cb}}_{\mu}$, defined in (2).
Lemma 4.
Fix $\mu\in(0,1)$ and consider the function $F_{\mu}:[0,1]\to[1,+\infty)$ defined as
[…]
Figure 1: Pictorial representation of the main step in the proof of Theorem 1. $1+\beta_{1}(x-\mu)$ dominates $E$ for $x\in[0,\mu)$, while $1+\beta_{0}(x-\mu)$ dominates $E$ for $x\in(\mu,1]$. If there is $\beta^{\star}\in(\beta_{0},\beta_{1})$, then we can find $u_{0}\in[0,\mu)$ and $u_{1}\in(\mu,1]$ such that $E(u_{0})>1+\beta^{\star}(u_{0}-\mu)$ and $E(u_{1})>1+\beta^{\star}(u_{1}-\mu)$, represented by the points $U_{0}$ and $U_{1}$ being above the purple line $1+\beta^{\star}(x-\mu)$. The probability measure $\hat{P}$ supported on $\{u_{0},u_{1}\}$ with mean $\mu$ is in $\mathcal{H}_{\mu}$. We have $\mathbb{E}_{\hat{P}}[E]>1$. Indeed, this expected value corresponds to the vertical coordinate of the point $U^{\star}$, the intersection of the line connecting $U_{0}$ and $U_{1}$ with the vertical line at $x=\mu$. This is a contradiction if $E$ is an e-variable, in which case it must be that $\beta_{0}\geq\beta_{1}$.
$F_{\mu}\;:\;x\mapsto\begin{cases}1+\frac{1}{\mu}(x-\mu)&\text{if $x\geq\mu$;}\\ 1+\frac{1}{\mu-1}(x-\mu)&\text{if $x<\mu$.}\end{cases}$
For any $x\in[0,1]$ there is $P_{x}\in\mathcal{H}_{\mu}$ such that $P_{x}(\{x\})=1/F_{\mu}(x)>0$. Moreover, $F_{\mu}\succeq E$ for every $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$.
Proof.
For $x\geq\mu$, let $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x})\delta_{0}$. Then, $P_{x}\in\mathcal{H}_{\mu}$ and $P_{x}(\{x\})=\mu/x=1/F_{\mu}(x)$. With a similar reasoning, if $x<\mu$ we can find a measure $P_{x}$ in $\mathcal{H}_{\mu}$, supported on $\{x,1\}$, with mass $1/F_{\mu}(x)$ on $x$.
Now, to check that $F_{\mu}$ majorises all the e-variables, fix $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$ and $x\in[0,1]$. Consider a $P_{x}\in\mathcal{H}_{\mu}$ with mass $1/F_{\mu}(x)$ on $x$. Then, $1\geq\mathbb{E}_{P_{x}}[E]\geq P_{x}(x)E(x)=E(x)/F_{\mu}(x)$, and we conclude. ∎
Theorem 1.
For any $\mu\in(0,1)$, the coin-betting e-class $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is the optimal e-class for $\mathcal{H}_{\mu}$.
Proof.
First, let us show that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class. For a pictorial representation of this part of the proof, the reader is invited to look at Figure 1. Fix an arbitrary $E\in\mathcal{E}_{\mathcal{H}_{\mu}}$. Define the sets
$B_{0}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in[0,\mu)}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\quad\text{and}\quad B_{1}=\left\{\beta\in I_{\mu}\;:\;\inf_{x\in(\mu,1]}\big{(}E_{\beta}(x)-E(x)\big{)}\geq 0\right\}\,,$
where we recall that $I_{\mu}=[(\mu-1)^{-1},\mu^{-1}]$ and $E_{\beta}:x\mapsto 1+\beta(x-\mu)$. Both sets are closed and convex (as they are intersections of closed and convex sets). By Lemma 4, $(\mu-1)^{-1}\in B_{0}$ and $\mu^{-1}\in B_{1}$, so $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu^{-1}]$, for some $\beta_{0}$ and $\beta_{1}$ in $I_{\mu}$. We will now show that $B_{0}\cap B_{1}\neq\varnothing$, or equivalently that $\beta_{0}\geq\beta_{1}$. Assume that this was not the case and $\beta_{0}<\beta_{1}$. Let $\beta^{\star}\in(\beta_{0},\beta_{1})$. Then, $\beta^{\star}\notin B_{0}$ and $\beta^{\star}\notin B_{1}$. In particular, there are $u_{0}\in[0,\mu)$ and $u_{1}\in(\mu,1]$ such that $E(u_{0})>E_{\beta^{\star}}(u_{0})$ and $E(u_{1})>E_{\beta^{\star}}(u_{1})$. As $\mu\in(u_{0},u_{1})$, there is $\hat{P}\in\mathcal{H}_{\mu}$ with support $\{u_{0},u_{1}\}$. Note that $\mathbb{E}_{\hat{P}}[E_{\beta^{\star}}]=1+\beta^{\star}(\mathbb{E}_{\hat{P}}[X]-\mu)=1$. But $E$ is strictly larger than $E_{\beta^{\star}}$ on $\mathrm{Supp}(\hat{P})$, and so $\mathbb{E}_{\hat{P}}[E]>1$, which is a contradiction since $E$ is an e-variable. We conclude that $\beta_{0}\geq\beta_{1}$ and so there exists $\hat{\beta}\in B_{0}\cap B_{1}$. By construction, $E_{\hat{\beta}}\in\mathcal{E}^{\mathrm{cb}}_{\mu}$ and $E_{\hat{\beta}}(x)\geq E(x)$ for all $x\in[0,1]$ different from $\mu$. For $x=\mu$, by Lemma 4 we have $E(\mu)\leq F_{\mu}(\mu)=1=E_{\hat{\beta}}(\mu)$, and so $E\preceq E_{\hat{\beta}}$. As the choice of $E$ was arbitrary, $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class.
Now that we have established that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class, by Lemma 3 we only need to show that all its elements are maximal to conclude that it is optimal. Fix $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$, and consider an e-variable $\hat{E}\in\mathcal{E}_{\mathcal{H}_{\mu}}$ such that $\hat{E}\succeq E$. Fix any $x\in[0,1]$. By Lemma 4, there is $P\in\mathcal{H}_{\mu}$ such that $P(\{x\})>0$. Since $E\in\mathcal{E}^{\mathrm{cb}}_{\mu}$, we have $\langle P,E\rangle=1$, and so $0\leq P(\{x\})(\hat{E}(x)-E(x))\leq\mathbb{E}_{P}[\hat{E}-E]=\mathbb{E}_{P}[\hat{E}]-1\leq 0$. Since $P(\{x\})>0$, we get $\hat{E}(x)=E(x)$ and, $x$ being arbitrary, $\hat{E}=E$. Hence $E$ is maximal, as it has no strict majoriser. ∎
5 Optimality beyond single-round e-variables
Up to this point, we have focused on hypotheses defined as subsets of the space $\mathcal{P}_{[0,1]}$, which cannot capture dependencies across multiple rounds. In such setting, the dependency structure among observations is fixed and assumed a priori, rather than being subject to testing. Yet, e-variable-based testing naturally extends to more general hypotheses that span multiple rounds and can account for sequential or dependent data structures. In the second part of the paper, we turn our attention to this richer framework. Remarkably, a slight adaptation of the proof strategy used for Theorem 1 and depicted in Figure 1 allows us to show that coin-betting-based testing is also optimal in the setting where the sequence of observations has a fixed conditional mean $\mu$. We will first show that this is the case with multi-round e-variables for a fixed time horizon, and then extend the result to testing with e-processes. Conversely, we will show that when testing the i.i.d. assumption, the coin-betting approach no longer yields the optimal e-class.
For technical reasons (primarily related to measurability, see discussion in Section 7), for the rest of this work we restrict ourselves to the case of a random variable taking values in a finite (known) set. Therefore, henceforth we let $\mathcal{X}\subseteq\mathbb{R}$ denote a finite set. For simplicity, we assume that $\min\mathcal{X}=0$ and $\max\mathcal{X}=1$, although all our results extend readily to any case where the minimum and maximum values are finite. We note that since $\mathcal{X}$ has finite cardinality, every function $\mathcal{X}^{t}\to\mathbb{R}$ is automatically Borel measurable.
5.1 Optimal e-class for multi-round e-variables
For any $\mu\in(0,1)$ and $T\geq 1$, we let
$\mathcal{H}_{\mu}^{T}=\big{\{}P\in\mathcal{P}_{\mathcal{X}^{T}}\,:\,\mathbb{E}_{P}[X_{t}|X^{t-1}]=\mu\,,\;\forall t\in[1:T]\big{\}}\,,$
where $\mathbb{E}_{P}[X_{1}|X^{0}]=\mathbb{E}_{P}[X_{1}]$. Note that $\mathcal{H}_{\mu}^{T}$ is a hypothesis on $\mathcal{X}^{T}$. As such, its set of e-variables will consist of Borel functions from $\mathcal{X}^{T}$ to $\mathbb{R}$. We denote the set of all e-values relative to $\mathcal{H}_{\mu}^{T}$ as $\mathcal{E}_{\mu}^{T}$.
In a coin-betting game, the wealth of the player at round $T$ is a non-negative function of the observations, which takes the form $M_{T}(x^{T})=\prod_{t=1}^{n}E_{t}(x_{t})$, where $E_{t}$ is in the form $E_{\lambda_{t}}$, with $\lambda_{t}$ chosen as a function of the past observations $x^{t-1}$. It is straightforward to check that under the hypothesis $\mathcal{H}_{\mu}^{T}$, the expectation of $M_{T}$ is always exactly one ($M_{T}$ is a martingale), which shows that $M_{T}$ is an e-variable. Next, we show that the e-class of e-variables in this form coincides is the optimal e-class for $\mathcal{H}_{\mu}^{T}$.
We let $\Lambda^{T}_{\mu}$ be a set of $T$-tuples of functions defined as
$\Lambda^{T}_{\mu}=\big{\{}\lambda^{T}=(\lambda_{1},\dots,\lambda_{T})\,:\,\text{$\lambda_{t}$ is a function from $\mathcal{X}^{t-1}$ to $I_{\mu}$}\big{\}}\,.$
Given $\lambda^{T}\in\Lambda^{T}_{\mu}$, we define the e-variable $E_{\lambda^{T}}\in\mathcal{E}^{T}_{\mu}$ as
$E_{\lambda^{T}}(x^{T})=\prod_{t=1}^{T}\big{(}1+\lambda_{t}(x^{t-1})(x_{t}-\mu)\big{)}\,.$
We define the T round coin-betting e-class $\mathcal{E}^{\mathrm{cb},T}_{\mu}=\{E_{\lambda^{T}}\,:\,\lambda^{T}\in\Lambda_{\mu}^{T}\}$.
Theorem 2.
Fix any $T\geq 1$ and $\mu\in(0,1)$. $\mathcal{E}^{\mathrm{cb},T}_{\mu}$ is the optimal e-class for $\mathcal{H}_{\mu}^{T}$.
The proof of Theorem 2 shares many similarity with that of Theorem 1. In particular, we use the same technique to establish the maximality of e-variables of the form $E_{\lambda^{T}}$. A geometric argument analogous to the one illustrated in Figure 1 is then employed, as part of an induction recursion, to prove that the e-class of interest is majorising. However, the detailed proof being somewhat lengthy, we defer it to the appendix.
5.2 A remark on the i.i.d. case
From what we have shown so far, in a $T$ round game coin-betting yields all and only the maximal e-variables, when testing the hypothesis that the conditional mean is some fixed $\mu\in(0,1)$. Yet, this is not any more the case if we consider a more restrictive hypotheses, stating that the observations are i.i.d. and with mean $\mu$. More concretely, let
^$\hat{\mathcal{H}}_{\mu}^{T}=\big{\{}P=Q^{\otimes T}\,:\,Q\in\mathcal{H}_{\mu}\big{\}}\,$
[…]
To give a concrete example of the mismatch between the e-variables for the conditional and independent scenarios, we consider the “highly symmetrical” case of $\mathcal{X}=\{0,1/2,1\}$, with $\mu=1/2$ and $T=2$. We show at the end appendix that the e-variables in this setting are the functions $E:\{0,1/2,0\}^{2}\to[0,+\infty)$ that satisfy
$\xi_{0}\leq 1\,;\qquad\xi_{1}\leq 1+\sqrt{(1-\xi_{0})(1-\xi_{2})}\,;\qquad\xi_{2}\leq 1\,,$
where $\xi_{0}=E(1/2,1/2)$, $\xi_{1}=(E(1,1/2)+E(1/2,1)+E(1/2,0)+E(0,1/2))/4$, and $\xi_{2}=(E(1,1)+E(1,0)+E(0,1)+E(0,0))/4$. Clearly, the optimal e-class here consists of those e-variables satisfying the constraint on $\xi_{1}$ with equality. We note that all the e-variables in $\mathcal{E}_{1/2}^{T}$ satisfy $\xi_{0}=\xi_{1}=\xi_{2}=1$. As expected, these are indeed maximal e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$, but there are maximal e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$ that are not in $\mathcal{E}_{1/2}^{T}$. The function that equals $4$ on $x^{2}=(1,1/2)$, and 00 everywhere else, is an example of a (maximal) e-variable for ^$\hat{\mathcal{H}}_{1/2}^{2}$, which is not an e-variable for $\mathcal{H}_{1/2}^{2}$.
We leave as a non-trivial open question to characterise the e-variables for ^$\hat{\mathcal{H}}^{T}_{\mu}$ in the general case.
6 E-processes for testing the fixed conditional mean
A key advantage of many testing procedures involving e-variables is anytime validity. One way to achieve this is through the round-wise testing-by-betting approach, which gives rise to a super-martingale (cf. Lemma 1). However, as we have already pointed out, this is not the most general approach. Rather than only focusing on super-martingales, one can instead use any non-negative process whose expectation is upper bounded by one, regardless the stopping time. These are known as e-processes, a generalization of super-martingales, which can enable powerful tests even in scenarios where martingale-based methods lack power (Ramdas et al., 2022b).
As in the previous section, here $\mathcal{X}$ denotes a finite subset of $[0,1]$, with $\min\mathcal{X}=0$ and $\max\mathcal{X}=1$. We denote as $\mathcal{X}^{\infty}$ the space of sequences $(x_{t})_{t\geq 1}\subseteq\mathcal{X}$. For convenience, we will sometimes write $x^{\infty}$ to mean $(x_{t})_{t\geq 1}$. We endow the set $\mathcal{X}^{\infty}$ with the product sigma-field generated by cylinder sets, and we denote as $\mathcal{P}_{\mathcal{X}^{\infty}}$ the space of probability measures on $\mathcal{X}^{\infty}$. For a fixed $\mu\in(0,1)$, in this section we will focus on the hypothesis class
$\mathcal{H}_{\mu}^{\infty}=\big{\{}P\in\mathcal{P}_{\mathcal{X}^{\infty}}\,:\,\mathbb{E}_{P}[X_{t}|X^{t-1}]=\mu\,,\;\forall t\geq 1\big{\}}\,.$
This time, we will not aim to study the e-variables for $\mathcal{H}^{\infty}_{\mu}$, as these would be functions that take a whole sequence as argument. Conversely, we are interested in tools that allow us to stop the test if we think enough data-points have been observed to decide whether or not the hypothesis has to be rejected. To make this rigorous we first need some definitions. First, we call a finite stopping time a function $\tau:\mathcal{X}^{\infty}\to\mathbb{N}$ such that, whenever for some $x^{\infty}$ and $y^{\infty}$ in $\mathcal{X}^{\infty}$ and some $t\in\mathbb{N}$ we have $\tau(x^{\infty})=\tau(y^{\infty})=t$, then $x^{t}=y^{t}$. [^8] It is straightforward to verify that in our setting this definition is equivalent to the perhaps more classical notion of finite e-process as a $\tau$ such that the sets $\{\tau\leq t\}$ are $\mathcal{F}_{t}$ measurable. We say that $\tau$ is bounded by $T\geq 0$ if its image is contained in $[0:T]$. We let $\mathcal{T}_{T}$ denote the set of all stopping times bounded by $T$.
Definition 6.
An e-process (for $\mathcal{H}_{\mu}^{\infty}$) is a sequence $(E_{t})_{t\geq 0}$, of non-negative functions $E_{t}:\mathcal{X}^{t}\to[0,+\infty)$, such that, for every $P\in\mathcal{H}^{\infty}_{\mu}$, any $T\geq 0$ and any $\tau\in\mathcal{T}_{T}$,
$\mathbb{E}_{P}[E_{\tau}]\leq 1\,.$
We denote as $\mathcal{E}_{\mu}^{\infty}$ the set of all the e-processes for $\mathcal{H}_{\mu}^{\infty}$. We call e-process class any subset of $\mathcal{E}_{\mu}^{\infty}$. Similarly to what we have done for the e-variables, we say that the e-process $E$ majorises the e-process $E^{\prime}$ (we write $E\succeq E^{\prime}$) if for every $t\geq 0$ we have $E_{t}\succeq E^{\prime}_{t}$. We say that an e-process class $\mathcal{E}$ is majorising if for every $E\in\mathcal{E}^{\infty}_{\mu}$, there is $E^{\prime}\in\mathcal{E}$ such that $E\preceq E^{\prime}$. We say that $\mathcal{E}$ is optimal if it is majorising and is contained in every other majorising e-process class. We call an e-process $E$ maximal if it is not strictly majorised by any other e-process. With the same reasoning as in Lemma 3, one can show that the optimal e-process class exists if, and only if, the class of all the maximal e-processes is majorising. In such case the optimal class is unique and coincides precisely with all and only the majorising e-processes.
We let $\Lambda_{\mu}^{\infty}$ be the set of all sequences $\lambda^{\infty}=(\lambda_{t})_{t\geq 1}$, such that $\lambda_{t}:\mathcal{X}^{t-1}\to I_{\mu}$. For any $\lambda^{\infty}\in\Lambda_{\mu}^{\infty}$, we define $E^{\lambda^{\infty}}=(E^{\lambda^{\infty}}_{T})_{T\geq 0}$ via
$E^{\lambda^{\infty}}_{T}(x^{T})=E_{\lambda^{T}}(x^{T})=\prod_{t=1}^{T}\big{(}1+\lambda_{t}(x^{t-1})(x_{t}-\mu)\big{)}\,,$
for $T\geq 1$, and $E_{0}=1$. These are the e-processes associated to the wealth of a coin-betting player. As $E^{\lambda^{\infty}}$ defines a martingale under $\mathcal{H}_{\mu}^{\infty}$, it is also an e-process for such hypothesis. We define the coin-betting e-process class $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}=\{E^{\lambda^{\infty}}\,:\,\lambda^{\infty}\in\Lambda^{\infty}_{\mu}\}$.
The next result shows that for $\mathcal{H}_{\mu}^{\infty}$ the coin betting approach is optimal also at the level of e-processes. The proof (reported in the appendix) always follows the ideas introduced in the proof of Theorem 1.
Theorem 3.
For any $\mu\in(0,1)$, $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is the optimal e-process class for $\mathcal{H}_{\mu}^{\infty}$.
7 Perspectives
Theorem 1 gives a rigorous sense to the claim that the coin-betting formulation is optimal among the e-variable-based approaches to build confidence sequences for $\mu^{\star}$. Theorem 2 and Theorem 3 extended this optimality result to testing the fixed conditional mean of a sequence of bounded observations. To formalise these claims, we introduced the notions of majorising and optimal e-classes, which, in the author’s view, are of independent interest in the context of sequential testing, beyond the scope of this paper. The main novelty of these concepts lies in defining “optimality” in terms of e-classes, rather than individual e-variables. This perspective contrasts with the notion of log-optimality of a single e-variable with respect to an alternative hypothesis, widely discussed in the literature (e.g., Koolen and Grünwald, 2022; Grünwald et al., 2024; Larsson et al., 2024).
The concept of maximality for e-variables, as introduced in this paper, is essentially equivalent to the classical statistical notion of admissibility. An e-variable is considered maximal if no other e-variable strictly dominates it. Likewise, the idea of a majorising e-class corresponds to the classical concept of a complete class, with the optimal e-class representing the minimal complete class of e-variables. For a detailed discussion of these ideas in traditional statistics we refer to Lehmann and Romano (2022). In this work, however, we adopt terminology from partially ordered set (poset) theory to highlight the fact that with e-variables these properties follows from the classical partial order of dominance among functions.
The concept of admissibility in the context of e-variables and e-processes was previously introduced by Ramdas et al. (2022a), whose definition is closely aligned with our notion of maximality. The key distinction lies in the type of dominance considered: their framework often relies on almost sure dominance, whereas our definition of maximality requires everywhere dominance. [^9] The two definitions becomes of course equivalent when $\mathcal{X}$ is finite, but they are not in the case of $\mathcal{X}=[0,1]$, which is the setting of our analysis on single-round e-variables. This stricter requirement is motivated by the will of developing a theory that remains valid even in the absence of a known alternative, where any value in $\mathcal{X}$ might occur at the next observation. Although Ramdas et al. (2022a) provide some necessary and sufficient conditions for admissibility, we emphasize that their results are insufficient to establish the optimality of the coin-betting e-class. Specifically, their approach assumes the existence of a reference measure for the hypothesis, [^10] A hypothesis $\mathcal{H}$ admits a ”reference measure” if there exists a Borel measure $Q$ such that $P\ll Q$ for all $P\in\mathcal{H}$. a condition not satisfied by $\mathcal{H}_{\mu}$. More precisely, their results can show that each individual e-variable in the coin-betting e-class is maximal, but not that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ forms a majorising e-class. This distinction, though subtle, is crucial: proving the optimality of coin-betting requires analyzing the collective properties of a set of e-variables (or equivalently, wealth processes), rather than evaluating them in isolation. It is precisely the fact that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class that guarantees nothing is lost by relying on the coin-betting approach. This point is subtle but essential: establishing the optimality of coin-betting requires studying the properties of a set of e-variables (or equivalently, wealth-processes) rather than focusing on single e-variables or wealth-processes individually. Crucially, it is precisely the fact that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is a majorising e-class that ensures that nothing is lost by relying on the coin-betting approach.
To the author’s knowledge, this is the first work to rigorously examine the optimality of the coin-betting formulation, as previous discussion on the topic has relied solely on heuristic arguments. For instance, Waudby-Smith and Ramdas (2023) justify restricting to the coin-betting e-class by noticing that $\mathcal{E}^{\mathrm{cb}}_{\mu}$ is precisely the set ¯$\bar{\mathcal{E}}_{\mathcal{H}_{\mu}}$ of e-variables whose expectation equals one under every $P\in\mathcal{H}_{\mu}$. However, this property does not generally imply the optimality of the e-class, as it merely indicates that all its elements are maximal. As a simple counterexample, consider $\mathcal{H}^{\prime}_{\mu}=\{P\in\mathcal{P}\,:\,\langle P,X\rangle\leq\mu\}$. Then, ¯$\bar{\mathcal{E}}_{\mathcal{H}^{\prime}_{\mu}}=\{1\}$, while the optimal e-class exists and consists of all functions in the form $E_{\beta}:x\mapsto 1+\beta(x-\mu)$, with $\beta\in[0,\mu^{-1}]$. Although one might argue that for each of the maximal e-variable there is at least one measure that bring the expectation to one (Grünwald, 2024 calls such property sharpness), sharpness alone is not enough to ensure maximality ($\mathbf{1}_{\mu}$ is a sharp e-variable for $\mathcal{H}_{\mu}$, but it is clearly not maximal).
Two papers (Clerico, 2024; Larsson et al., 2025) have appeared after the first preprint of this work, both extending the results on characterising single-round e-variables for hypotheses defined via linear constraints, a framework that includes the testing for the mean as a special case. Their analyses, however, rely on considerably more abstract and technically advanced proof techniques. Conversely, a key strength of the present paper is to provide a simple and neat argument (illustrated in Figure 1) that works well in the simple setting considered. Interestingly, this same argument is also at the base of the proofs of Theorem 2 and Theorem 3. We also remark that a similar strategy was adopted in the proof of Lemma 2 in Wang (2025). That said, for the multi-round setting, the approach does give rise to non-trivial measurability issues when $\mathcal{X}$ is uncountable. These difficulties already emerge in the two-round case: although one can show that every e-variable must be dominated by a product of the form $(1+\lambda_{1}x_{1})(1+\lambda_{2}(x_{1})x_{2})$, this is constructed in a way that gives no guarantee on the measurability of $\lambda_{2}$. Whether this issue can be resolved remains an open question. Nonetheless, even in the simplified setting we adopt by requiring $\mathcal{X}$ to have finite cardinality [^11] A countable $\mathcal{X}$ would also suffice to solve the measurability issue, though the final step of the proof of Theorem 3 would require more delicate handling., this work is, to the best of the author’s knowledge, the first to fully characterise all e-processes for the natural problem of testing a fixed conditional mean.
To conclude, let us remark that this paper does not aim to delve into the design of effective coin-betting strategies for obtaining tight confidence sequences. For this, we refer the interested reader to the thorough analysis and discussion by Orabona and Jun (2023) and Waudby-Smith and Ramdas (2023). Finally, although we have focused on random variables taking values in $[0,1]$ for clarity of exposition, all the results presented readily extend to any bounded closed interval.
Acknowledgements
The author would like to thank Gergely Neu, Peter D. Grünwald, Nishant A. Mehta, Hamish E. Flynn, and Claudia M. Chanu, for the insightful discussions that inspired this work. I would also like to thank Nick W. Koning and Aditya Ramdas for their constructive feedback, which helped improve this work following its initial appearance, and Wouter M. Koolen for helpful discussions on e-processes. The author acknowledges the use of GPT-4o during the redaction of this paper, in order to polish the presentation. All AI-generated text was subsequently reviewed and edited by the author, who takes full responsibility for the content of this manuscript. This project was funded by the European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation programme (grant agreement 950180).
Appendix
Proof of Theorem 2
Proof.
We start by showing that the class of functions considered is an e-class, and that all its elements are maximal e-variables. So, fix any $\lambda^{T}_{\mu}\in\Lambda^{T}$. As all the $\lambda_{t}$ are valued in $I_{\mu}$, it is clear that $E_{\lambda^{T}}$ is non-negative. Moreover, for any $P\in\mathcal{H}_{\mu}^{T}$ it is easily checked that $\mathbb{E}_{P}[E_{\lambda^{T}}]=1$, which ensures that $E_{\lambda^{T}}$ is an e-variable. For the maximality we can use the same argument as in the proof of Theorem 1. Let $\hat{E}\in\mathcal{E}^{T}_{\mu}$ be such that $\hat{E}\succeq E_{\lambda^{T}}$. Fix any $x^{T}\in\mathcal{X}^{T}$. By Lemma 4, for each $t$ we can find a $P_{t}\in\mathcal{H}^{1}_{\mu}$ such that $P_{t}(\{x_{t}\})>0$. Let $P=P_{1}\otimes\dots\otimes P_{T}$. Then we have that $P(\{x^{T}\})>0$. In particular, we have that $0\leq P(\{x^{T}\})(\hat{E}(x^{T})-E_{\lambda^{T}}(x^{T}))\leq\mathbb{E}_{P}[\hat{E}-E_{\lambda^{T}}]=\mathbb{E}_{P}[\hat{E}]-1\leq 0$. Since $P(\{x^{T}\})>0$, $\hat{E}(x^{T})=E_{\lambda^{T}}(x^{T})$, and we conclude that $E_{\lambda^{T}}$ is maximal.
To show that the e-class that we consider is majorising we proceed by induction on $T$. From Theorem 1 we know that the statement is true for $T=1$. For $T\geq 2$. let us assume that the claim holds for $T-1$, and let us show that it is also true for $T$. Let $E$ be an e-variable for $\mathcal{H}^{T}_{\mu}$. First, we show that, for any $Q\in\mathcal{H}^{T-1}_{\mu}$ and any $x\in\mathcal{X}$, we have
$\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq F_{\mu}(x)\,,$
with $F_{\mu}$ as in Lemma 4. Fix $Q$ and $x\geq\mu$ in $\mathcal{X}$. Define $V_{x}=P_{x}\otimes Q$, where $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x})\delta_{0}$. Clearly, $V_{x}\in\mathcal{H}_{\mu}^{T}$. We have that $1\geq\mathbb{E}_{V_{x}}[E]=\tfrac{\mu}{x}\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]+\big{(}1-\tfrac{\mu}{x}\big{)}\mathbb{E}_{Q}[E(0,X_{2},\dots,X_{T})]$, and so $\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq x/\mu=F_{\mu}(x)$. The case for $x\leq\mu$ is essentially analogous. In particular, the following two sets are non-empty:
$B_{0}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap[0,\mu)\,,\,\forall Q\in\mathcal{H}_{\mu}^{T-1}\,,\,\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq 1+\beta(x-\mu)\big{\}}\,;$
$B_{1}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap(\mu,1]\,,\,\forall Q\in\mathcal{H}_{\mu}^{T-1}\,,\,\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq 1+\beta(x-\mu)\big{\}}\,.$
$B_{0}$ and $B_{1}$ are intersections of closed intervals (one per each allowed $x$ and $Q$), and as such they must be closed intervals. We can find $\beta_{0}$ and $\beta_{1}$ such that $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu^{-1}]$. We will show that $\beta_{0}\geq\beta_{1}$, and hence that $B_{0}\cap B_{1}\neq\varnothing$. The argument is essentially the same used in the proof of Theorem 1 (cf. also Figure 1). Assume that $\beta_{0}<\beta_{1}$, and let $\beta^{\star}\in(\beta_{0},\beta_{1})$. As $\beta^{\star}\notin B_{0}$, there is $Q_{0}\in\mathcal{H}^{T-1}_{\mu}$ and $u_{0}<\mu$ in $\mathcal{X}$ such that $\mathbb{E}_{Q_{0}}[E(u_{0},X_{2},\dots,X_{T})]>1+\beta^{\star}(u_{0}-\mu)$. Similarly, there must be $Q_{1}\in\mathcal{H}_{\mu}^{T-1}$ and $u_{1}>\mu$ in $\mathcal{X}$, such that $\mathbb{E}_{Q_{1}}[E(u_{1},X_{2},\dots,X_{T})]>1+\beta^{\star}(u_{1}-\mu)$, since $\beta^{\star}\notin B_{1}$. Let $V=\frac{u_{1}-\mu}{u_{1}-u_{0}}\delta_{u_{0}}\otimes Q_{0}+\frac{\mu-u_{0}}{u_{1}-u_{0}}\delta_{u_{1}}\otimes Q_{1}$. Then, $V\in\mathcal{H}^{T}_{\mu}$. Moreover,
$\mathbb{E}_{V}[E]>\tfrac{u_{1}-\mu}{u_{1}-u_{0}}\big{(}1+\beta^{\star}(u_{0}-\mu)\big{)}+\tfrac{\mu-u_{0}}{u_{1}-u_{0}}\big{(}1+\beta^{\star}(u_{1}-\mu)\big{)}=1\,,$
which is a contradiction since $E\in\mathcal{E}^{T}_{\mu}$. We have thus established that $B_{0}\cap B_{1}$ is non-empty, and in particular there is $\lambda_{1}\in I_{\mu}$ such that, for every $x\in\mathcal{X}$ and every $Q\in\mathcal{H}_{\mu}^{T-1}$,
$\mathbb{E}_{Q}[E(x,X_{2},\dots,X_{T})]\leq 1+\lambda_{1}(x-\mu)\,.$
Fix $y\in\mathcal{X}$ such that $1+\lambda_{1}(y-\mu)\neq 0$. Define $E_{y}:\mathcal{X}^{T-1}\to\mathbb{R}$ via $E_{y}(x^{2:T})=E(y,x^{2:T})/(1+\lambda_{1}(y-\mu))$, where we recall that $x^{2:T}$ is a shorthand for $(x_{2},\dots,x_{T})$. $E_{y}$ is non-negative, since by definition of $I_{\mu}$ we know that $1+\lambda(y-\mu)$ is non-negative for any $\lambda\in I_{\mu}$. From what we have already shown, $E_{y}$ is an e-variable for $\mathcal{H}^{T-1}_{\mu}$. Thanks to the inductive hypothesis, for each $t\in[2:T]$ there exist a function $\tilde{\lambda}_{t}^{y}:\mathcal{X}^{t-2}\to I_{\mu}$ such that $E_{y}(x^{2:T})\leq\prod_{t=2}^{T}(1+\tilde{\lambda}_{t}^{y}(x^{2:(t-1)}))(x_{t}-\mu))$, for any $x^{2:T}\in\mathcal{X}^{T-1}$. For all $t\in[2:T]$, and every $x^{t-1}\in\mathcal{X}^{t-1}$, define
$\lambda_{t}(x^{t-1})=\begin{cases}0&\text{if $1+\lambda_{1}(x_{1}-\mu)=0$};\\ \big{(}1+\lambda_{1}(x_{1}-\mu)\big{)}\tilde{\lambda}_{t}^{x_{1}}(x^{2:(t-1)})&\text{otherwise.}\end{cases}$
So far, we have proved that for any $x^{T}\in\mathcal{X}^{T}$, such that $1+\lambda_{1}(x_{1}-\mu)\neq 0$, we have that $E(x^{T})\leq E_{\lambda^{T}}(x^{T})$. We only need to check that this is still true when $01+\lambda_{1}(x_{1}-\mu)=$ 0. In such case, $E$ is $Q$-almost surely null for all $Q\in\mathcal{H}^{T-1}_{\mu}$, which implies that $E(x^{T})=0$, and so we conclude. ∎
Proof of Theorem 3
Proof.
We start by recalling and introducing some definitions. For $T\geq 1$, $\Lambda^{T}_{\mu}$ is the set whose elements are $T$-tuple $\lambda^{T}=(\lambda_{1},\dots,\lambda_{T})$, with $\lambda^{t}:\mathcal{X}^{t-1}\to I_{\mu}$. We endow $\Lambda^{T}_{\mu}$ with the product topology of point-wise convergence. Specifically, a sequence $(\lambda^{t})_{t\geq 0}\to\tilde{\lambda}^{t}$ if, and only if, $\lambda_{t}\to\tilde{\lambda}_{t}$ point-wise for all $t\leq T$. For convenience, we will henceforth refer to this topology as the $\Lambda$-topology. It is straightforward to check that under the $\Lambda$-topology $\Lambda_{\mu}^{T}$ is compact, since $I_{\mu}$ is bounded and $\mathcal{X}$ has finite cardinality. Our proof strategy will rely on an inductive reasoning. For this, it is convenient to introduce the notion of a $T$-e-process, for $T\geq 1$. This is a sequence $(E_{t})_{t\geq 0}$ of non-negative functions $E_{t}:\mathcal{X}^{t}\to[0,+\infty)$, such that $\mathbb{E}_{P}[E_{\tau}]\leq 1$ for any $\tau\in\mathcal{T}_{T}$, and any $P\in\mathcal{H}_{\mu}^{\infty}$. We remark that an e-process is a $T$-e-process for all $T\geq 1$.
We are now ready to start the proof. First, we will show that $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is a majorising e-process class. For any $T\geq 1$ and any $T$-e-process $E$, let $\tilde{\Lambda}^{T}_{E,\mu}\subseteq\Lambda^{T}_{\mu}$ denote the set of the $\lambda^{T}$ such that $E_{t}\preceq E_{\lambda^{t}}$, for all $t\in[1:T]$. We note that this is a closed and compact set under the $\Lambda$-topology that we introduced earlier. We now show that for every $T\geq 1$ and every $T$-e-process $E$, the set $\tilde{\Lambda}^{T}_{E,\mu}$ is non-empty. We prove this claim by induction. Since $E_{1}$ must be an e-variable for $\mathcal{H}_{\mu}^{1}$, as a consequence of Theorem 1, $\tilde{\Lambda}^{1}_{E,\mu}$ is non-empty. Now, fix $T\geq 2$ and a $T$-e-process $E$, and assume that, for any $(T-1)$-e-process $E^{\prime}$, $\tilde{\Lambda}^{T-1}_{E^{\prime},\mu}$ is non-empty. We will show that this implies that $\tilde{\Lambda}_{E,\mu}^{T}$ is non-empty. Fix an arbitrary $x\in\mathcal{X}$ and define the process $E^{x}=(E^{x}_{t})_{t\geq 1}$ such that $E_{t}^{x}(x^{t})=E_{t+1}(x,x_{1},\dots,x_{t})$, for any $x^{t}\in\mathcal{X}^{t}$. Define the sets
$B_{0}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap[0,\mu),\forall Q\in\mathcal{H}^{\infty}_{\mu},\forall\tau\in\mathcal{T}_{T-1},\,\mathbb{E}_{Q}[E^{x}_{\tau}]\leq 1+\beta(x-\mu)\big{\}}\,;$
$B_{1}=\big{\{}\beta\in I_{\mu}\,:\,\forall x\in\mathcal{X}\cap(\mu,1],\forall Q\in\mathcal{H}^{\infty}_{\mu},\forall\tau\in\mathcal{T}_{T-1},\,\mathbb{E}_{Q}[E^{x}_{\tau}]\leq 1+\beta(x-\mu)\big{\}}\,.$
We claim that $B_{0}$ and $B_{1}$ are non-empty. Indeed, we now show that $\mu^{-1}\in B_{1}$. Analogously one can prove that $(\mu-1)^{-1}\in B_{0}$. The argument here follows closely the one depicted in Figure 1. Fix $x\in\mathcal{X}\cap[0,\mu)$, $Q\in\mathcal{H}_{\mu}^{\infty}$, and $\tau\in\mathcal{T}_{T-1}$. Define $\tau_{x}:\mathcal{X}^{\infty}\to\mathbb{N}$ as $\tau_{x}(x^{\infty})=1+\tau(x^{2:\infty})$, if $x_{1}=x$, otherwise $\tau_{x}(x^{\infty})=1$. It is easily checked that $\tau\in\mathcal{T}_{T}$. Also, we let $P_{x}=\frac{\mu}{x}\delta_{x}+(1-\frac{\mu}{x}\delta_{0})$, and we define $V_{x}=P_{x}\otimes Q$ [^12] Note that both $V_{x}$ and $Q$ are measures on $\mathcal{P}_{\mathcal{X}^{\infty}}$. Here we mean that $V_{x}(\{x^{\infty}\})=P_{x}(\{x_{1}\})Q(\{x^{2:\infty}\})$ for all $x^{\infty}\in\mathcal{X}^{\infty}$. Clearly, $V_{x}\in\mathcal{H}_{\mu}^{\infty}$. Since $E$ is a $T$-e-process, we have
$1\geq\mathbb{E}_{V_{x}}[E_{\tau_{x}}]=(1-\tfrac{\mu}{x})E_{1}(0)+\tfrac{\mu}{x}\mathbb{E}_{Q}[E^{x}_{\tau}]\geq\tfrac{\mu}{x}\mathbb{E}_{Q}[E^{x}_{\tau}]\,.$
So, $E_{Q}[E^{x}_{\tau}]\leq\frac{x}{\mu}=1+\mu^{-1}(x-\mu)$, and so $\mu^{-1}\in B_{1}$.
Since both $B_{0}$ and $B_{1}$ can be written as intersections of non-empty closed intervals, we can find $\beta_{0}$ and $\beta_{1}$ such that $B_{0}=[(\mu-1)^{-1},\beta_{0}]$ and $B_{1}=[\beta_{1},\mu^{-1}]$. As usual, we want to show that $B_{0}\cap B_{1}$ is non-empty, or equivalently that $\beta_{1}\leq\beta_{0}$. Assume that this was not the case and there is $\beta^{\star}\in(\beta_{0},\beta_{1})$. Since $\beta^{\star}\notin B_{0}$, there must be $u_{0}<\mu$ in $\mathcal{X}$, $Q_{0}\in\mathcal{H}_{\mu}^{\infty}$, and $\tau_{0}\in\mathcal{T}_{T-1}$ such that $\mathbb{E}_{Q_{0}}[E^{u_{0}}_{\tau_{0}}]>1+\beta^{\star}(u_{0}-\mu)$. Similarly, as $\beta^{\star}\notin B_{1}$, one can find $u_{1}>\mu$ in $\mathcal{X}$, $Q_{1}\in\mathcal{H}_{\mu}^{\infty}$, and $\tau_{1}\in\mathcal{T}_{T-1}$, such that $\mathbb{E}_{Q_{1}}[E^{u_{1}}_{\tau_{1}}]>1+\beta^{\star}(u_{1}-\mu)$. Define $\tau:\mathcal{X}^{\infty}\to\mathcal{N}$ via $\tau(x^{\infty})=1+\tau_{0}(x^{2:\infty})$, if $x_{1}<\mu$, $\tau(x^{\infty})=1+\tau_{1}(x^{2:\infty})$ if $x_{1}>\mu$, and $\tau(x^{\infty})=1$ if $x_{1}=\mu$. Also, let $V=\frac{u_{1}-\mu}{u_{1}-u_{0}}\delta_{u_{0}}\otimes Q_{0}+\frac{\mu-u_{0}}{u_{1}-u_{0}}\delta_{u_{1}}\otimes Q_{1}$. Then, we have that
$\mathbb{E}_{V}[E_{\tau}]=\frac{u_{1}-\mu}{u_{1}-u_{0}}\mathbb{E}_{Q_{0}}[E^{u_{0}}_{\tau_{1}}(X_{2},\dots,X_{T})]+\frac{\mu-u_{0}}{u_{1}-u_{0}}\mathbb{E}_{Q_{1}}[E^{u_{1}}_{\tau_{1}}(X_{2},\dots,X_{T})]>1\,.$
However, this is a contradiction since $\tau\in\mathcal{T}_{T}$, $V\in\mathcal{H}_{\mu}^{\infty}$, and $E$ is a $T$-e-process. We can thus conclude that $B_{0}\cap B_{1}\neq\varnothing$ and, in particular, there must be $\lambda_{1}\in I_{\mu}$ such that, for every $x\in\mathcal{X}$, every $Q\in\mathcal{H}_{\mu}^{\infty}$, and every $\tau\in\mathcal{T}_{T-1}$, we have
$\mathbb{E}_{Q}[E^{x}_{\tau}]\leq 1+\lambda_{1}(x-\mu)\,.$
Now, for any arbitrary $x\in\mathcal{X}$ such that $1+\lambda_{1}(x-\mu)\neq 0$. Define the process $\tilde{E}^{(x)}=(\tilde{E}^{(x)}_{t})_{t\geq 0}$ via
$\tilde{E}^{(x)}_{t}(x^{t})=\frac{E_{t+1}(x,x_{1},\dots,x_{t})}{1+\lambda_{1}(x-\mu)}\,.$
Let us show that $\tilde{E}^{(x)}$ is a $(T-1)$-e-process. It is clearly non-negative, as $\lambda_{1}\in I_{\mu}$ implies $1+\lambda_{1}(x-\mu)\geq 0$. Fix $Q\in\mathcal{H}_{\mu}^{\infty}$ and $\tau\in\mathcal{T}_{T-1}$. From what we have shown, we have
$\mathbb{E}_{Q}[\tilde{E}^{(x)}_{\tau}]=\frac{\mathbb{E}_{Q}[E^{x}_{\tau}]}{1+\lambda_{1}(x-\mu)}\leq 1\,.$
In particular, the inductive argument yields the existence of $\tilde{\lambda}_{(x)}^{T-1}\in\Lambda^{T-1}_{\mu}$, such that for all $t\in[1:T-1]$ we have $\tilde{E}^{(x)}\preceq E_{\tilde{\lambda}^{t}_{(x)}}$. Let $\lambda^{T}\in\Lambda^{T}_{\mu}$ be defined via $\lambda_{t}(x^{T})=\big{(}1+\lambda_{1}(x_{1}-\mu)\big{)}\tilde{\lambda}_{(x_{1}),t}(x_{2},\dots,x_{T})$, where we set $\tilde{\lambda}_{(x),t}$ identically equal to zero whenever $01+\lambda_{1}(x-\mu)=$ 0. We will now show that $\lambda^{T}\in\tilde{\Lambda}_{E,\mu}^{T}$. Fix any $x^{T}\in\mathcal{X}^{T}$. If $1+\lambda_{1}(x_{1}-\mu)\neq 0$, then for any $t\in[1:T]$ we have that
$E_{t}(x^{t})=\big{(}1+\lambda_{1}(x_{t}-\mu)\big{)}E^{(x_{1})}_{t-1}(x_{2},\dots,x_{t})\leq\prod_{s=1}^{t}\big{(}1+\lambda_{s}(x^{s-1})(x_{s}-\mu)\big{)}=E_{\lambda^{t}}(x^{t})\,.$
If $01+\lambda_{1}(x_{1}-\mu)=$ 0, then, for every $t\in[1:T-1]$, $\mathbb{E}_{Q}[E_{t}^{x_{1}}]=0$ for every $Q\in\mathcal{H}^{\infty}$, which implies that $E_{t}^{x_{1}}$ is identically null. In particular, $E_{t}(x^{t})=0=E_{\lambda^{t}}(x^{t})$. This shows that $\lambda^{T}\in\tilde{\Lambda}_{E,\mu}^{T}$, and so $\tilde{\Lambda}_{E,\mu}^{T}$ is non-empty.
What we have shown so far allows to say that given an e-process $E$, for every $T\geq 1$ we can find a $\lambda^{T}\in\Lambda^{T}$ such that $E_{t}\leq E_{\lambda^{t}}$ for every $t\in[1:T]$. However, this does not tell us yet that there is a $\lambda^{\infty}\in\Lambda^{\infty}_{\mu}$ such that the same property holds. To show this, we leverage the compactness of $\tilde{\Lambda}^{T}_{E,\mu}$ under the $\Lambda$-topology. For $T^{\prime}\geq T$, let $\tilde{\Lambda}^{T|T^{\prime}}_{E,\mu}$ denote the set of $\lambda^{T}\in\Lambda^{T}$ such that there is $\tilde{\lambda}^{T^{\prime}}\in\tilde{\Lambda}_{E,\mu}^{T^{\prime}}$ satisfying, for all $t\in[1:T]$, $\lambda_{t}=\tilde{\lambda}_{t}$. It is easily checked that $\tilde{\Lambda}^{T|T^{\prime}}_{E,\mu}$ is compact in the $\Lambda$-topology. Moreover, if $T^{\prime\prime}\geq T^{\prime}$, then $\tilde{\Lambda}^{T|T^{\prime\prime}}_{E,\mu}\subseteq\tilde{\Lambda}^{T|T^{\prime}}_{E,\mu}$. Since the intersection of countably many non-empty compact sets is non-empty, we can define $\tilde{\Lambda}^{1|\infty}_{E,\mu}=\bigcap_{T\geq 1}\tilde{\Lambda}^{1|T}_{E,\mu}$, which is a non-empty set. Now, fix $\lambda_{1}\in\tilde{\Lambda}^{1|\infty}_{E,\mu}$. We have that $E_{1}\preceq E_{\lambda_{1}}$. By construction $\lambda_{1}\in\tilde{\Lambda}^{1|2}_{E,\mu}$, and so we can find a $\lambda_{2}$ such that $(\lambda_{1},\lambda_{2})\in\tilde{\Lambda}^{2}_{E,\mu}$. We can go on like this indefinitely. In particular, we can construct a sequence $\lambda^{\infty}$ such that $\lambda^{T}\in\tilde{\Lambda}^{T}_{E,\mu}$ for all $T\geq 1$. This concludes the proof that $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is a majorising e-process class.
We are left with checking that every e-process in $\mathcal{E}^{\mathrm{cb},\infty}_{\mu}$ is maximal. Fix $\lambda^{\infty}\in\Lambda^{\infty}_{\mu}$, and let $E$ be an e-process such that $E\succeq E^{\lambda^{\infty}}$. Clearly, we must have $E_{0}=1=E_{0}^{\lambda^{\infty}}$. Now, fix $T\geq 1$. We have that $E_{T}\succeq E_{\lambda^{T}}$. Since $E$ and $E^{\lambda^{\infty}}$ are e-processes, choosing the stopping time that is identically equal to $T$ shows that $E_{T}$ and $E_{\lambda^{T}}$ are $T$-round e-variables. In particular, by Theorem 2, $E_{\lambda^{T}}$ is maximal, and so $E_{T}\succeq E_{\lambda^{T}}$ implies that $E_{T}$ and $E_{\lambda^{T}}$ coincide. Since this is true for all $T\geq 1$, we conclude. ∎
Characterising the e-variables for ^$\hat{\mathcal{H}}_{1/2}^{2}$ with $\mathcal{X}=\{0,1/2,1\}$
Let $\mathcal{X}=\{0,1/2,1\}$, and $\mu=1/2$. Let $Q\in\mathcal{H}_{1/2}^{1}$ and let $q=Q(\{1/2\})$. The mean constraints reads $Q(\{1\})+q/2=1/2$, which implies that $Q(\{1\})=(1-q)/2$. The normalisation yields $Q(\{0\})=1-q-(1-q)/2=(1-q)/2$. This shows that there is a one-to-one correspondence between $[0,1]$ and $\mathcal{H}_{1/2}^{1}$. Now, since by definition there is a one-to-one correspondence between $\mathcal{H}_{1/2}^{1}$ and ^$\hat{\mathcal{H}}_{1/2}^{2}$, we obtain that we can parametrise ^$\hat{\mathcal{H}}_{1/2}^{2}$ by $[0,1]$. For $q\in[0,1]$, we hence let $P_{q}$ be the (unique) element of $\mathcal{H}_{1/2}^{2}$ that gives mass $q^{2}$ to $(1/2,1/2)$.
Let $E:\mathcal{X}^{2}\to[0,+\infty)$ be a non-negative function. We have that
$\mathbb{E}_{P_{q}}[E]=q^{2}\xi_{0}+2q(1-q)\xi_{1}+(1-q)^{2}\xi_{2}=(\xi_{0}+\xi_{2}-2\xi_{1})q^{2}-2(\xi_{0}-\xi_{1})q+\xi_{2}\,,$
where $\xi_{0}=E(1/2,1/2)$, $\xi_{1}=(E(1,1/2)+E(1/2,1)+E(1/2,0)+E(0,1/2))/4$, and $\xi_{2}=(E(1,1)+E(1,0)+E(0,1)+E(0,0))/4$. Now, $E$ is in ^$\hat{\mathcal{E}}_{1/2}^{2}$ if, and only if,
$\max_{q\in[0,1]}\big{(}(\xi_{0}+\xi_{2}-2\xi_{1})q^{2}-2(\xi_{0}-\xi_{1})q+\xi_{2}\big{)}\leq 1\,.$
In particular, checking for $q=0$ and $q=1$ implies that $\xi_{0}\leq 1$ and $\xi_{2}\leq 1$. If both $\xi_{0}$ and $\xi_{2}$ are equal to $1$, then the constraint on $\xi_{1}$ becomes $\max_{q\in[0,1]}(1-\xi_{1})(q^{2}-q)\leq 0$, which implies $\xi_{1}\leq 1$. We are left to check if when one among $\xi_{0}$ and $\xi_{2}$ is strictly less then $1$ a $\xi_{1}$ can be larger than one. In such case, we note that we have $\xi_{2}-\xi_{1}<0$ and $\xi_{0}-\xi_{1}<0$. This implies that the parabola is concave and achieve its maximum at $q^{\star}=(\xi_{0}-\xi_{1})/(\xi_{0}-\xi_{1}+\xi_{2}-\xi_{1})\in(0,1)$. The maximum is equal to $\xi_{2}(\xi_{1}-\xi_{0})^{2}/(2\xi_{1}-\xi_{0}-\xi_{2})$. Asking that this quantity is less than one reduces to the constraint $\xi_{1}\leq 1+\sqrt{(1-\xi_{0})(1-\xi_{2})}$.